### v1

# Deep Learning based Visually Rich Document Content Understanding: A Survey

Abstract.
Visually Rich Documents (VRDs) are essential in academia, finance, medical fields, and marketing due to their multimodal information content. Traditional methods for extracting information from VRDs depend on expert knowledge and manual labor, making them costly and inefficient. The advent of deep learning has revolutionized this process, introducing models that leverage multimodal information—vision, text, and layout—along with pretraining tasks to develop comprehensive document representations. These models have achieved state-of-the-art performance across various downstream tasks, significantly enhancing the efficiency and accuracy of information extraction from VRDs. In response to the growing demands and rapid developments in Visually Rich Document Understanding (VRDU), this paper provides a comprehensive review of deep learning-based VRDU frameworks. We systematically survey and analyze existing methods and benchmark datasets, categorizing them based on adopted strategies and downstream tasks. Furthermore, we compare different techniques used in VRDU models, focusing on feature representation and fusion, model architecture, and pretraining methods, while highlighting their strengths, limitations, and appropriate scenarios. Finally, we identify emerging trends and challenges in VRDU, offering insights into future research directions and practical applications. This survey aims to provide a thorough understanding of VRDU advancements, benefiting both academic and industrial sectors.
Visually Rich Document Understanding, Key Information Extraction, Question Answering, Entity Linking, Multimodal
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Information systems Information extraction
1. Introduction
1.1. Backgrounds
Visually Rich Documents (VRDs) are documents that contain a diverse mix of visual and textual elements designed to convey information in a comprehensive and visually engaging manner, which are ubiquitous in our daily lives and appear in domains such as finance, medicine, and academia. These documents integrate various types of visually rich elements, including paragraphs, tables, charts, diagrams, and photos. Both textual elements (such as paragraphs, lists, and captions) and visually rich elements (such as tables and figures), collectively referred to as document semantic entities, are essential for illustrating, explaining, and summarizing information. Common formats for VRDs include PDF (Portable Document Format), DOC/DOCX (Microsoft Word Document), and image files (JPEG, PNG). These documents are typically structured in a semi-structured or even unstructured manner, making their understanding and information extraction challenging.
The VRD Understanding (VRDU) field is dedicated to comprehending the structure of VRDs and extracting pertinent information, thus transforming unstructured or semi-structured content into a machine-readable format. VRDU mainly encompasses two tasks: Key Information Extraction (KIE) and Question Answering (QA). Key Information Extraction (KIE) aims to identify and extract values based on predefined keys. Depending on the model used, KIE can be approached as an entity retrieval task (KIE Task Formulation 1) or a sequence tagging task (Task Formulation 2), as shown in Figure 1. Visually Rich Document Question Answering (VRD-QA) involves answering questions in natural language based on contextual information in VRDs. As explained in Figure 1, the model must locate the answer within the document to respond to the input question. Beyond extractive methods, KIE and VRD-QA can also be formulated as generative tasks, where the model generates the desired value or answer auto-regressively based on the query or question and the document images. In addition to the above-mentioned tasks, another content understanding task is Entity Linking (EL), which is proposed to identify the semantic correlations between document entities (parent-child relations).
To address visually rich document understanding tasks, heuristic methods (Watanabe et al., 1995; Seki et al., 2007; Rusinol et al., 2013) and statistical machine learning techniques (Oliveira and Viana, 2017) have been used in domain-specific document applications, which require expert customization and are challenging to update. Recent advances in deep learning have introduced promising alternatives. LSTM and CNN-based models (Katti et al., 2018; Denk and Reisswig, 2019; Zhao et al., 2019), feature-driven approaches (Yu et al., 2021; Zhang et al., 2020; Wang et al., 2021a), and layout-aware pretrained frameworks (Xu et al., 2020a; Wang et al., 2022b; Hong et al., 2021), along with visual-integrated pretrained frameworks (Huang et al., 2022), have significantly improved document representation and achieved state-of-the-art performance in various downstream tasks. However, these models primarily focus on fine-grained features, such as grids and word/word pieces (textual tokens), and often struggle with computational complexity and the ability to capture global logical and layout correlations. In contrast, coarse-grained models, which operate at the entity level, address some limitations of fine-grained models but may omit detailed information, resulting in less comprehensive representations. To bridge these gaps, joint-grained frameworks (Li et al., 2021c; Yu et al., 2022; Bai et al., 2022) and LLM-based frameworks (Luo et al., 2024; Liu et al., 2024b) have been developed.
With the rapid development and increasing demands in Visually Rich Document Understanding (VRDU), various model architectures, multimodal learning methods, and pretraining techniques have been introduced to consistently enhance the performance of specific or multiple VRDU tasks. A comprehensive review of VRDU frameworks based on deep learning is provided, systematically surveying and analyzing existing methods and benchmark datasets, which are categorized based on adopted strategies and downstream tasks. Additionally, different techniques used in VRDU models are compared, focusing on feature representation and fusion, model architecture, and pretraining methods, with their strengths, limitations, and appropriate scenarios highlighted. Finally, emerging trends and challenges in visually rich document content understanding are identified, offering insights into future research directions and practical applications.
1.2. Scope
This paper is devoted to surveying and analysing existing work on understanding deep learning-based document content. Papers on the following topics will be reviewed and summarized in this paper.
  • The paper focuses on document content understanding tasks that will be included in this paper, mainly including document key information extraction, visual question answering and entity linking. This paper will not summarise the models and datasets that only focus on document structure understanding tasks.
  • A paper proposed a novel deep-learning VRDU model for addressing one or more Visually Rich Document Content Understanding (VRD-CU) tasks. Section 2.2.1 will briefly describe typical document content understanding approaches in a heuristically or traditional machine learning way.
  • This paper compiles widely used VRDU benchmark datasets and recently proposed VRDU datasets from top-tier conferences and journals since 2019. All included datasets are publicly available.
  • This paper focuses on the multimodal capabilities for achieving a holistic understanding of entire, multi-page documents, in contrast to previous document understanding surveys that primarily focus on isolated subtasks.
It should be noted that this article only investigates models and datasets designed for the entire document or the document page rather than the specific document components such as table, chart. Therefore, models and datasets proposed for table detection, table structure recognition, and chart or plot question answering will not be summarized in the main body of this survey.
1.3. Related Surveys
Several surveys offer a comprehensive overview of general document understanding tasks. These surveys primarily focus on document layout analysis (Subramani et al., 2020), table extraction (Liu et al., 2023), named entity recognition (Ehrmann et al., 2023), and document image analysis (Lombardi and Marinai, 2020) across diverse document types, including invoices (Saout et al., 2024) and historical document (Ehrmann et al., 2023; Lombardi and Marinai, 2020). In particular, computer vision-based research focuses on scanned document analysis and structure understanding. While these studies have advanced document image analysis, they often focus on fragmented subtasks and fall short of providing a holistic understanding of an entire document on multiple pages.
Recent advancements in deep learning have fueled the emergence of VRDU tasks, which demand complex document content understanding capabilities. These tasks include key information extraction, question answering, and document entity linking. However, existing surveys have not adequately addressed the unique challenges and opportunities presented by VRDU, with a specific emphasis on deep learning-based multimodal approaches (Cui et al., 2021). To bridge this gap, this survey aims to provide a comprehensive overview of VRDU frameworks and datasets, including multimodal feature extractions and fusions in both mono and multi-task VRD models.
1.4. Contributions
The main contributions of this paper can be summarized as follows:
  • The paper provides a detailed review and systematic categorization of VRDU frameworks and benchmark datasets, organized based on adopted strategies and downstream tasks.
  • It critically examines and compares different techniques used in VRDU models, focusing on feature representation and fusion, model architecture, and pretraining methods, highlighting their strengths, limitations, and appropriate scenarios.
  • The paper identifies emerging trends and challenges in visually rich document content understanding, offering insights into future research directions and practical applications.
1.5. Survey Structure
Section 1 in this survey discusses the background of visually rich document understanding and emphasizes the aim, scope, and contributions of this paper. Section 2 provides additional background knowledge, including definitions of the VRDU tasks covered in this research and the development of document understanding. Sections 3 and 3 review significant frameworks for mono-task and multi-task VRDU, respectively. Section 5 compiles benchmark datasets for three VRDU subtasks: scanned receipts, forms, and single and multiple-page documents from diverse domains. Section 6 critically examines and compares the techniques adopted by various models, highlighting their strengths and limitations. Finally, Section 7 summarizes the trends and future directions in this field.
2. Background
2.1. Task Definition
Based on the differences in the purpose and application scenarios, visually rich document understanding tasks are classified into three categories, including Key Information Extraction, Question Answering and Entity Linking.
Figure 1. Visually rich document content understanding task clarifications.
  • Key Information Extraction: refers to identifying and extracting the relevant information based on the given text queries. Distinct pre-defined queries can be defined based on the domain of targeting documents and practical demands. For example, the crucial information of scanned receipts contains ”Store Name”, ”Address”, ”Item” and ”Price”, while for the financial reports, ”Company Name”, ”Share Holder Name”, ”Number of Interests” may be the critical information need to be extracted.
  • Question Answering: is a task of answering questions about a VRD by using natural languages. Based on the answer types, it can be divided into extractive VQA and abstractive VQA. The answer from extractive VQA is directly extracted from the target document, while abstractive VQA requires generative answers based on comprehensively understanding questions and related VRDs.
  • Document Entity Linking: refers to identifying the semantic relations between document entities to construct the logical structure of the input document image.
2.2. Development of Document Understanding
2.2.1. Traditional Approaches
Rule-based methods, as highlighted in several studies (Watanabe et al., 1995; O’Gorman, 1993; Seki et al., 2007; Rusinol et al., 2013), have demonstrated high precision in domain-specific applications. However, these methods have several drawbacks: they are manually intensive, costly, and require expert intervention for tailored customization. Additionally, they are inflexible, often necessitating frequent manual updates, even for minor modifications. In response to these limitations, machine learning-based approaches have been proposed for document understanding. For instance, SVM-based methods have been utilized for layout understanding (Oliveira and Viana, 2017), and TF-IDF techniques combined with hand-crafted features have been applied for extracting information from invoices. Furthermore, rule-based and statistical models have been used for entity extraction. Despite these advancements, machine learning methods rely heavily on human intervention and domain-specific expertise. They are time-consuming and often deliver suboptimal performance. Moreover, the majority of these methods typically depend on single-modality data, restricting them to either layout, text, or visual inputs.
2.2.2. Single Modality-based Approaches
With the advancement of deep learning, deeper model architectures such as CNNs (Katti et al., 2018; Yang et al., 2017) have emerged, and pretrained language (Devlin, 2018; Liu et al., 2019b) and vision models (Ren et al., 2015; He et al., 2017) are now commonly used as robust baselines for understanding VRD content and structure. Due to the multimodal nature of VRDs, which involves the integration of text, vision, and layout, researchers are increasingly focusing on leveraging this combined information to achieve significant improvements in various downstream VRDU tasks.
2.2.3. Cross Modality based Approaches
Considering the multimodal nature of VRDs, many frameworks propose various methods to encode multimodal information, including text, vision, and layout, and fuse them effectively. Text and vision information is normally encoded by pretrained backbones such as BERT (Devlin, 2018) or RoBERTa (Liu et al., 2019b) for textual features, Faster-RCNN (Ren et al., 2015) or Mask-RCNN (He et al., 2017) for visual features. For layout information, different encoding methods are introduced, including linear projection (Wang et al., 2020b), 2D positional encoding (Xu et al., 2020a), and attention bias to allow the proposed models to be layout-sensitive. Different feature fusion methods are introduced including summing up (Yu et al., 2021), concatenation (Lee et al., 2022b), attention-based contextual learning (Majumder et al., 2020a), and prompting (He et al., 2023). However, most of those frameworks leverage implicit knowledge from pretrained backbones with a task-orientated shadow adapter for specific VRDU downstream tasks such as KIE (Lee et al., 2023; Wang and Shang, 2022; Chen et al., 2023; Cao et al., 2023b) or EL (Zhang et al., 2021; Hu et al., 2023; Carbonell et al., 2021). Those frameworks tend to achieve delicate performance on specific tasks or document formats instead of acquiring a generalised model to represent documents comprehensively.
2.2.4. Multimodal Pre-training Approches
Inspired by the success of BERT-style models (Devlin, 2018; Liu et al., 2019b) in acquiring knowledge through self-supervised learning, pretrained document understanding models have emerged to harness self-supervised or supervised pretraining tasks from extensive document collections. LayoutLM (Xu et al., 2020a), the first encoder-only VRDU model, utilizes self-supervised tasks, such as masked vision-language modelling, with text and layout information. Subsequent models have expanded on this by integrating layout information (Wang et al., 2022b; Tu et al., 2023; Li et al., 2021c) and visual cues (Xu et al., 2020b; Huang et al., 2022) through multimodal transformers. While encoder-only models have shown significant improvements on various benchmark datasets (Jaume et al., 2019; Park et al., 2019; Huang et al., 2019; Mathew et al., 2021; Harley et al., 2015), they often require detailed annotations and are limited by fixed input lengths. To address these limitations, encoder-decoder frameworks (Tang et al., 2022; Kim et al., 2022; Davis et al., 2022) and prompt-based methods for LLMs/MLLMs (Luo et al., 2024; Liu et al., 2024b) have been developed, enhancing layout awareness and performance in VRDU tasks. However, a significant gap remains in effectively applying these models in real-world scenarios with zero shot.
3. Mono-Task Document Understanding Frameworks
Visually rich document content understanding encompasses several independent downstream tasks tailored to different application scenarios and user demands. This section introduces methods focused on specific document understanding tasks. Many of these models aim to provide comprehensive document representation and integrate target-oriented modules or techniques to improve performance and efficiency across various VRDU tasks. The models for three VRDU downstream tasks—Key Information Extraction (KIE), Entity Linking (EL), and Visual Question Answering (VQA)—are introduced and summarized with insights into current trends.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=3cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Mono-task based models (Content Understanding), for tree=fill=lime!60 [Key Information Extraction, for tree=fill=cyan!40 [Feature-driven Models, for tree=fill=cyan!25 [Chargrid(Katti et al., 2018); CUTIE(Zhao et al., 2019);BERTgrid(Denk and Reisswig, 2019); ACP(Palm et al., 2019);XYLayoutLM(Gu et al., 2022), for tree=fill=cyan!10] ] [Joint-learning Frameworks, for tree=fill=cyan!25 [TRIE(Zhang et al., 2020); VIES(Wang et al., 2021a), for tree=fill=cyan!10] ] [Relation-aware Models, for tree=fill=cyan!25 [Majumder et al. (Majumder et al., 2020a); Liu et al. (Liu et al., 2019a); PICK(Yu et al., 2021); FormNet(Lee et al., 2022b); FormNetv2(Lee et al., 2023), for tree=fill=cyan!10] ] [Few/Zero-shot Learning Models, for tree=fill=cyan!25 [LASER(Wang and Shang, 2022); Chen et al.(Chen et al., 2023); Cheng et al.(Cheng et al., 2020); QueryForm(Wang et al., 2023a), for tree=fill=cyan!10] ] [Prompt-based Frameworks, for tree=fill=cyan!25 [GenKIE(Cao et al., 2023b); ICL-D3IE(He et al., 2023); LMDX(Perot et al., 2023), for tree=fill=cyan!10] ] ] [Entity Linking, for tree=fill=blue!40 [Entity-level Linking, for tree=fill=blue!20 [DocStruct(Wang et al., 2020b); Zhang et al.(Zhang et al., 2021); KVPFormer.(Hu et al., 2023), for tree=fill=blue!10] ] [Token-level Linking, for tree=fill=blue!20 [Carbonell et al. (Carbonell et al., 2021); SPADE (Hwang et al., 2021);DocTR(Liao et al., 2023), for tree=fill=blue!10] ] ] [Visual Question Answering, for tree=fill=teal!40 [Single Page Frameworks, for tree=fill=teal!25 [Please refer Section 3. , for tree=fill=teal!10] ] [Multi-Page Frameworks, for tree=fill=teal!25 [Hi-VT5(Tito et al., 2023); GRAM(Blau et al., 2024); Kang et al. (Kang et al., 2024), for tree=fill=teal!10] ] ] ]
Figure 2. Mono-task visually rich document understanding models
3.1. Key Information Extraction
Key Information Extraction (KIE), a typical natural language processing task, refers to the task of identifying and extracting crucial pieces of information from textual data. Unlike typical name entity recognition methods for addressing plain text, VRDs contain visually rich entities like tables and charts, as well as spatial and logical layout arrangements to enhance the challenge of extracting crucial information. Although plain-text pretrained language models, such as BERT (Devlin, 2018), RoBERTa (Liu et al., 2019b), and ALBERT (Lan et al., 2019), are widely used as solid baselines on many benchmark datasets, more recent works introduce layout-aware pretrained models such as LayoutLM families (Xu et al., 2020a, b; Huang et al., 2022), LiLT (Wang et al., 2022b), Bros (Hong et al., 2021) to enhance the document representation by leveraging visual and layout information and achieving SoTA performance on several downstream tasks (see Section 3). This section will mainly focus on models specifically proposed for the document KIE models or only evaluated on KIE benchmark datasets including FUNSD (Jaume et al., 2019), CORD (Park et al., 2019), SROIE (Huang et al., 2019), XFUND (Xu et al., 2021), etc. Based on innovative aspects, we categorise KIE frameworks into five types: Feature-driven models use multimodal cues for rich feature representation. Joint Learning frameworks integrate auxiliary tasks to enhance document representation. Relation-aware models leverage spatial or logical relations via graphs or masked attention mechanisms. Few/Zero-shot learning frameworks explore methods for extracting key information with minimal labelled data, often using transfer learning. Prompt-based frameworks use structured prompts to guide specific information extraction from pretrained models or LLM/MLLMs. These categories represent diverse approaches to improving Key Information Extraction (KIE) by leveraging specific model designs and learning strategies tailored to document understanding challenges.
3.1.1. Feature-driven Models
In the initial phase, certain Recurrent Neural Network (RNN)-based models (Lample et al., 2016) were introduced, primarily focusing on key information extraction tasks from plain text. However, these approaches overlook the importance of visual cues and layout information. Hence, several multimodal frameworks with feature-driven designs have been proposed to generate more representative document representations.
Chargrid (Katti et al., 2018) first mentioned the significance of 2D structure for document KIE and designed a character-box-based chargrid to convert the textual and 2D layout structure into coloured visual cues feed into CNN. Instead of using fine-grained character information, CUTIE (Zhao et al., 2019) and BERTgrid (Denk and Reisswig, 2019) utilise various word embedding methods with bounding box coordinates to persevere the layout structure. ACP (Palm et al., 2019) leverages attention mechanism and multi-aspect features, including visual, semantic (both character and word) and spatial, into dilated CNN for capturing both short and long-term dependencies of each word piece.
Joint learning KIE frameworks are proposed to leverage multi-level features to mitigate the information gap between various focused tasks. TRIE (Zhang et al., 2020) firstly offers an end-to-end framework for simultaneously conducting Optical Character Recognition (OCR) and KIE. OCR module will generate multi-aspect features, including positional, visual and textual aspects. Adaptively trainable weighting mechanisms are adopted to generate the fused embedding, followed by a Bi-LSTM-based Entity Extraction Module to conduct the final prediction. VIES (Wang et al., 2021a) is introduced to use the multi-level cues of vision, position and text to generate more comprehensive representations. Both token and segment (entity) level positional and visual features are acquired from text detection modules, and dual-level textual features are gathered by text recognition brunch, fused by a self-attention-based fusion module for sequence labelling.
3.1.2. Relation-aware Models
Instead of only leveraging multimodal information, leveraging the inherent spatial or logical relations between multi-level components may lead to more robust and comprehensive document representations. Majumder et al. (Majumder et al., 2020a) propose a model which starts by generating candidates for each field using type-specific detectors and key phrases. The neural model scores these candidates by learning dense representations that consider both textual content and spatial positioning, allowing for accurate information extraction across different document templates by focusing on the candidates’ relevance to the fields. Graph-based frameworks are increasingly favoured for modelling document elements’ spatial and logical relationships. This trend involves meticulously defining distinct graph structures and employing graph convolution techniques to incorporate these relationships into feature representations seamlessly.
Liu et al. (Liu et al., 2019a) first utilise a graph convolution module on top of the BiLSTM-CRF for key information extraction. Each document is a fully connected graph where the node is the textual representation $T$ of the document entity $E_{i}\in\mathbb{E}$, and the edge is related to the spatial relation between $E_{i}$ and another entity $E_{j}$. It is defined as $e_{ij}=[x_{ij},y_{ij},\frac{w_{i}}{h_{i}},\frac{h_{j}}{h_{i}},\frac{w_{j}}{h_{i}}]$, where $x_{ij}$ and $y_{ij}$ are horizontal and vertical distance between $E_{i}$ and $E_{j}$. The self-attention-based graph convolution is performed on node-edge-node triplets (concatenated nodes and edge embeddings) to learn contextually with all other nodes. PICK (Yu et al., 2021) adopts a similar way to construct the document graph but utilising multimodal node representations, including transformer-encoded textual embedding and CNN-encoded visual embedding, as well as the edge embedding is updated to $e_{ij}=[x_{ij},y_{ij},\frac{w_{i}}{h_{i}},\frac{h_{j}}{h_{i}},\frac{w_{j}}{h_{i}},\frac{S_{j}}{S_{i}}]$ where $S$ is the sentence length of corresponding entity. Additionally, to get the node embedding for the downstream tagging task, a soft adjacent matrix-based graph learning layer is applied to obtain the task-specific node representations.
FormNet (Lee et al., 2022b) introduces an end-to-end framework with a new attention mechanism, Rich Attention, and graph framework to make an order/distance sensitive long sequence transformer. In Rich Attention, the model introduces two order-based and pixel distance scores along the x/y axis. These are summed up with usual self-attention scores to enable the model order/distance awareness. Additionally, a Graph is applied to contextually learn the neighbouring token embedding before serialising the token sequence feeding into the transformer encoders. The graph nodes are text embedding of tokens, and the edges are relative positions between nodes. This could mitigate the imperfect serialization issue, enabling token representation capturing in more contexts. FormNetv2 (Lee et al., 2023) integrates image modality as the additional Graph edge feature to capture more visual cues  [^1] Please refer to Section 6.1.2 to get more details. To use contrastive loss to learn the multimodal graph representation, they first perform stochastic graph corruption to sample topology-corrupted and feature-corrupted graphs. Topology corruption randomly removes edges in the original graph, while feature corruption drops all modality features from nodes and edges. Then, the contrastive objective is to maximize agreement between a pair of tokens between corrupted and original graphs under a standard normalized temperature-scaled cross-entropy (Sohn, 2016) loss.
3.1.3. Few-shot Learning Frameworks
Extracting key information from VRDs based on deep learning frameworks typically involves more manual work to annotate the training data. However, acquiring large-scale, high-quality annotations in urgent and labouring-limited application scenarios is challenging. Thus, few-shot or one-shot frameworks emerged to extract the key information with less labouring annotations.
LASER (Wang and Shang, 2022) leverage the architecture from LayoutReader (Wang et al., 2021b) to reformulate the entity recognition task to sequence-labelling to generative model by embedding the entity type information (named label surface name) into the target sequence to enable the model label semantic-aware. LASER depends on a ”partially triangle” mask to use one encoder to encode source text and generate the prediction. The $n$ source text sequence, $\{t_{1},...,t_{n}\}$, is the text sequence by summing the relative word, spatial, and positional embedding, feeding into the $\mathcal{E}$ with full self-attention. The target sequence only attends previous tokens. The generative formulation is defined as:
(1) $t_{i-1},[B],t_{i},...,t_{j},[E],e_{1},...,e_{k},[T],t_{j+1}$
where $[B]$ and $[E]$ denote start and end of entity, $e_{i}$ are the label surface name; $[T]$ denotes the end of the label surface name. As label surface names and functional tokens do not exist, the learnable embeddings are applied to them to ensure the generative-progress layout-aware. A binary classification module is applied to classify the next generated token type of current generated token hidden states: from source or not, effectively controlling the generative series.
Chen et al. (Chen et al., 2023) introduce a novel framework for entity-level, N-way soft-K-shot extracting key information from VRDs, focusing on the challenges of extracting rare or unseen entities from documents with few examples. It leverages a meta-learning approach (Snell et al., 2017; Oreshkin et al., 2018), utilizing a hierarchical decoder and contrastive learning (ContrastProtoNet) for task personalization and improved adaptation to new entity types. Additionally, it introduces FewVEX, a dataset tailored for entity-level few-shot VDER, to facilitate research and benchmarking in this area. The framework’s effectiveness is demonstrated through significant improvements in robustness and performance over existing meta-learning baselines.
In the domain of one-shot scenarios, Cheng et al. (Cheng et al., 2020) presents a novel application of graphs. They utilize attention mechanisms to seamlessly transfer spatial relationships between static text regions ( key/landmark) and dynamic text regions (value/field) from support documents to query documents. This transfer enables the acquisition of label probability distributions for files. Additionally, a self-attention-based module is integrated to exploit relationships between field entities, facilitating the derivation of label transition probability distributions. Finally, belief propagation is harnessed for inference within a pairwise Conditional Random Field (CRF), resulting in an assessable end-to-end trainable pipeline.
3.1.4. Prompt-learning Frameworks
QueryForm (Wang et al., 2023a) introduce a query-based framework for zero-shot document key information extraction. A dual prompting-based mechanism, entity query (E-prompt) and schema query (S-prompt), is introduced for pre-training and transferring knowledge from large-scale weakly annotated pre-trained webpages to target domains. The pretraining target is highly aligned with fine-tuning to ensure the proposed framework can consistently make query-conditional predictions at both stages. The HTML tags acquire the E-prompt ($\bm{e_{p}}$) during pretraining, and the S-prompt ($\bm{\tilde{s}_{p}}$) is generated from webpage domains, while during the fine-tuning, $\bm{e_{p}}$ is predefined and S-prompt $\bm{s_{p}}$ is learnable vectors. Supposing $\bm{t}$ is the serialised text of the input document, the pre-training and fine-tuning target $\bm{\hat{y}}$ and $\bm{y}$ can be represented as:
(2) $\bm{\hat{y}}=\mathcal{F}([\bm{s_{p}};\mathcal{E}[\bm{e_{p}},\bm{x}]]),$
(3) $\bm{y}=\mathcal{F}([\mathcal{E}[\bm{\tilde{s}_{p}};\bm{e_{p}},\bm{x}]])$
where $\mathcal{E}$ and $\mathcal{F}$ in here are the feature encoder and the rest of the language model, respectively. The training objective is to minimise the cross-entropy loss between $\bm{\hat{y}}$ and $\bm{y}$.
Prompt learning is a technique in natural language processing where models are guided by specific prompts to produce or interpret targeted responses. With the rise of large-scale models, prompt learning can leverage contextual representation and implicit knowledge to enhance performance on targeted tasks. Since most LLMs (OpenAI, 2023; Touvron et al., 2023) and MLLMs (Liu et al., 2024a) are trained on plain text or natural images, layout-aware prompting and in-context learning (Brown et al., 2020) methods have been introduced to improve understanding in visually rich documents (VRD)
GenKIE (Cao et al., 2023b) proposes an encoder-decoder-based multimodal KIE framework to leverage prompt to adapt various datasets and better leverage multimodal information. Following (Xu et al., 2020a, b), different encoding methods are adopted to acquire textual, layout and visual embeddings. Byte Pair Encoding is used as language pretrained backbones, and the OCR extracted document content is concatenated with predefined prompts split by the ”[SEP]” token between OCR tokens and each prompt. The 2D-positional encoding introduced by (Xu et al., 2020a) is used to acquire the layout embedding of each OCR extracted token, and ResNet (He et al., 2016) is used to extract the visual representations following (Wang et al., 2022c). The multimodal representations are fed into an encoder to learn interactively between modalities. Prompts, inserted at the end of the encoder’s textual inputs, are either template-style or question-style. For the entity extraction task, the prompt specifies the target entity type, and the decoder generates the entity value (e.g., for ”Company is?”, the decoder outputs the company name). For the entity-labeling task, the prompt includes the value, and the decoder provides the entity type (e.g., for ”Es Kopi Rupa is [SEP]”, the decoder identifies the entity type).
ICL-D3IE (He et al., 2023) is the first framework to employ LLMs with in-context learning to extract key information from VRDs using the iterative updated diverse demonstrates. Before designing the initial diverse demonstrations, the most similar $n$ training documents to the $n$ test samples need to be selected by calculating the cosine-similarity of document representations encoded by Sentence-BERT (Reimers and Gurevych, 2019). Then, different types of demonstrations are introduced to integrate multiple-view context information into LLMs. Hard Demonstrations highlight the most challenging cases, are initially designed based on the incorrectly predicted cases from GPT-3 (Brown et al., 2020) predictions and are updated based on prediction results during the training process. Layout-aware demonstrations are created by selecting the adjacent hard segments to understand the positional relations. Formatting demonstrations are designed to guide LLMs in formatting the outputs for easy post-processing.
LMDX (Perot et al., 2023) designs a pipeline to use arbitrary LLMs to extract singular, repeated and hierarchical entities from VRDs. The document images are first fed into off-the-shelf OCR tools and are divided into smaller document chunks to be processed by the LLMs with accessible input length. Then, prompts are generated under XML-like tags to control the LLM’s responses and mitigate hallucination. Document Representation is a prompt contains the chunk content with the coordinates of OCR lines to bring layout modality to LLMs, where a text segment extracted by OCR tools is represented ¡text¿ [$x_{centre}$, $y_{centre}$]. After that, the task description and scheme representation prompts are designed to explain the task to accomplish and determine the output format. During inference, $N$ prompts with $K$ LLMs completions are generated to sample the correct answers.
3.1.5. Summary of Key Information Extraction
Several models like Chargrid (Katti et al., 2018) and ACP (Palm et al., 2019) have enhanced VRDU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilized by (Zhang et al., 2020; Wang et al., 2021a), aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialized models, which produce less representative features compared to those generated by large-scale pre-trained models like LayoutLM (Xu et al., 2020a) and SelfDoc (Li et al., 2021b). Documents typically exhibit specific layouts and logical structures, which has prompted many models (Yu et al., 2021; Lee et al., 2022b, 2023) to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot (Wang and Shang, 2022) and zero-shot (Wang et al., 2023a) frameworks, which leverage contrastive learning and innovative attention mechanisms. Additionally, prompt learning has been applied to distill implicit knowledge from large-scale layout-aware pre-trained models (Hong et al., 2021; Tu et al., 2023) and large language models (LLMs/MLLMs) (He et al., 2023; Perot et al., 2023). Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRDU optimization.
3.2. Document Entity Linking
Documents are normally structured hierarchically, where parent-child relations always exist in various documents, such as key-value pairs in forms and section paragraphs in reports or papers. Unlike most VRD key information extraction models, which focus on recognising the semantic entity categories in a sequence tagging task ignoring the relation between entities, linking the logical associations between document semantic entity pairs has recently been of greater interest.
3.2.1. Entity-Level Entity Linking
Document entity linking aims to identify the relation between document entities. Some frameworks use the known entity bounding boxes, ignoring the entity recognition step and mainly focusing on exploring the relation between input document entities.
DocStruct (Wang et al., 2020b) is the first entity linking framework by leveraging multimodal features of known document entities to predict the hierarchical structure between them. [CLS] token from a BERT-like pretrained language model is extracted for entity-level textual representation $T_{e}$, and an RNN is applied to encode the sequential RoI visual cues $V_{e}$ along the width of the ResNet-50 feature maps. Then, $T_{e}$ is concatenated with layout feature $P_{e}$, which is a linear projected vector of entity bounding box coordinates, $[x1,y1,x2,y2,x3,y3,x4,y4]$. The final entity representation $E$ can be represented:
(4) $\alpha=Sigmoid(W[T_{e},P_{e},V_{e}]+b),$
(5) $E=[T_{e};P_{e}]+\alpha V_{e},$
where $\alpha$ is used as a gate to control the influence of visual cues $V_{e}$ following (Wang et al., 2019). Then, for pair of entities $E_{i}$ and $E_{j}$, the probability of an existing parent-child relationship is represented as $P_{i\rightarrow j}=E_{i}ME_{j}$ where M is the matrix of asymmetric parameters to ensure the asymmetric relation between $E_{i}$ and $E_{j}$. Negative sampling (Mikolov et al., 2013b) is applied during training to handle data sparsity and balance problems.
Zhang et al. (Zhang et al., 2021) propose an entity-linking framework, SERA (Semantic Entity Relation extraction As dependency parsing) by formulating the entity relation prediction as a dependency parsing problem. LayoutLM (Xu et al., 2020a) is used to obtain textual representations at the entity level $T_{e}$ that are concatenated with the embedding of the linear projected entity label $L_{e}$ to obtain the final entity representation $E$, which could be formulated as $E=[T_{e},L_{e}]$. Various encoders, such as Vanilla Transformer, BiLSTM, and Graph, are adopted to learn the following contextual representations. Then, a Biaffine parser is adopted to calculate the score between $E_{i}$ and $E_{j}$. The biaffine parser score $p_{b}$ can be calculated by:
(6) $h_{i}^{key}=\sigma(W^{key}E_{i}+b^{key}),$
(7) $h_{j}^{value}=\sigma(W^{value}E_{j}+b^{value}),$
(8) $p_{b}=h_{i}^{key}W_{b1}h_{j}^{value}+h_{i}^{key}W_{b2}.$
where $h_{i}^{key}$ and $h_{j}^{value}$ are the hidden states after linear projection. To leverage the layout information, a two-dim layout vector $l_{ij}$ is used to get the layout feature score $p_{l}=W_{l}l_{ij}+b_{l}$, where $l_{ij}$ is the minimum distance between two entities along with width and height direction respectively. Then, the final score $p=p_{b}+p_{l}$ is used to calculate loss using binary or multi-label classification.
KVPFormer (Hu et al., 2023) formulate the entity linking to a question-answer problem by introducing a transformer-based encoder-decoder structure to leverage joint-grained information (token and entity) to predict the entity association. The textual representation $T$ of each document entity, $E$, is encoded by averaging the inner-token embedding from pretrained document understanding models, concatenating with the linear projected entity label embedding $l$, represented as $E=[T;l]$. Then, the entity representations will be fed into a Transformer encoder, which has a spatial compatibility attention bias into vanilla self-attention mechanism  [^2] Please refer to . A binary classifier determines which entity is a key entity or not. All detected key entities will be treated as question representation and fed into a DETR-style (Carion et al., 2020) decoder to predict corresponding answers in parallel. For each input question (key query), Top K answer candidates will be selected based on the $Sigmoid$ score and fed into a softmax layer to get the final prediction, named a coarse-to-fine answer prediction.
3.2.2. Token-level Entity Linking
As acquiring entity information needs prior knowledge from manual annotation or layout analysing models, some works utilise serialised OCR-extracted tokens as inputs to extract the structured relations. However, as the logical relation links semantic entities, token-level frameworks must group tokens into entities before exploring their association.
Carbonell et al. (Carbonell et al., 2021) introduces a framework comprising three modules for sequentially conducting token grouping, entity labelling and relation prediction. Firstly, each text token $t$ is represented by $[L_{t};T_{t}]$ where $L_{t}=[x,y,w,h]$ is the coco format bounding box coordinates of $t$ and $T_{t}$ is the work/representations. All tokens are fed into a token grouping GNN, $\mathcal{G}_{group}$, as a node where the edge between nodes is determined by k-NN to avoid high consumption of fully connected GNN. The $\mathcal{G}_{group}$ is trained on a link prediction task to predict the edge score between two nodes to group words, of which scores larger than a predefined threshold $\rho$. Then, the grouped words are fed into a Graph Attention Network (GAT) to use multi-head attention to aggregate words into entities and follow an MLP to conduct node classification to predict the category of each document entity. At last, another link prediction GNN, $\mathcal{G}_{link}$, is trained on edge classification based on aggregated entities.
SPADE (Hwang et al., 2021) formulate the token-level entity linking as a spatial dependency parsing task for serializing (ordering and grouping) tokens and predicting inter-group relation between grouped tokens. Firstly, a spatial text encoder is designed to make spatial-aware attention by introducing a relative spatial vector considering relative, physical and angle aspects. During this task, two binary matrices must be predicted $M_{g}$ for token grouping and inter-group linking $M_{l}$. The vertices comprised by a entity types $\mathbb{V}$ and sequence of tokens $\mathbb{T}$ the encoded entity category and token are represented as $c$ and $t$, respectively, and the relation score between vertices $v_{i}\rightarrow v_{j}$ can be calculated by:
(9) $h_{i}=\begin{cases}c_{v_{i}},&\text{for }v_{i}\in\mathbb{V},\\ W_{h}t_{v_{i}},&\text{otherwise}\end{cases},$
(10) $d=W_{d}v_{j},$
(11) $s_{0}=h_{i}^{T}W_{0}d,\ s_{1}=h_{i}^{T}W_{1}d.$
The probability is acquired by $p_{ij}=\frac{\exp(s_{0,j})}{\exp(s_{0,j})+\exp(s_{1,j})}$. An adjustable threshold is set to construct the $M_{g}$ or $M_{l}$.
DocTR (Liao et al., 2023) formulate the entity linking as an anchor word-based entity detection and association problem. Each document entity is represented by the anchor word to convert the entity extraction and linking to token-level tasks. It contains a Deformable DETR-based vision encoder to extract multi-scale visual feature extraction. A LayoutLM based language encoder is applied to encode token-level textual representations. The outputs from vision/language encoders are fed into the vision-language decoder with the language-conditional queries to conduct entity extraction and linking. The decoder queries are one-to-one mapping with language encoder inputs. The entity extraction task aims to predict whether the query underlying token level input is an anchor word and corresponding categories, while entity linking is acquired by
3.2.3. Summary
Several models like Chargrid (Katti et al., 2018) and ACP (Palm et al., 2019) have enhanced VRDU by integrating visual and textual information. Additionally, auxiliary tasks such as OCR, utilised by (Zhang et al., 2020; Wang et al., 2021a), aid in improving multimodal feature representations through joint training. However, these frameworks often rely on smaller, randomly initialised models, which produce less representative features compared to those generated by large-scale pretrained models like LayoutLM (Xu et al., 2020a) and SelfDoc (Li et al., 2021b). Documents typically exhibit specific layouts and logical structures, which has prompted many models (Yu et al., 2021; Lee et al., 2022b, 2023) to adopt graph-based approaches. These methods capture spatial and logical correlations among document elements, such as key-value pairs, leading to a more comprehensive document representation. While these frameworks have achieved improvements in document representation, their effectiveness hinges on having sufficient well-annotated training samples, which are time-consuming to acquire. This limitation has escalated the demand for few-shot (Wang and Shang, 2022) and zero-shot (Wang et al., 2023a) frameworks, which leverage contrastive learning and innovative attention mechanisms. Furthermore, prompt learning has been applied to distil implicit knowledge from large-scale layout-aware pretrained models (Hong et al., 2021; Tu et al., 2023) and large language models (LLMs/MLLMs) (He et al., 2023; Perot et al., 2023). Despite these advances, a performance gap remains between well-fine-tuned models and few/zero-shot frameworks, highlighting the ongoing challenges in VRDU optimization.
3.3. VRD Question Answering
Unlike key information extraction, which targets specific details within document images, answering natural language questions involves interpreting more complex intentions and requires models to facilitate interactive understanding between the queries and document representations (Ding et al., 2022). The introduction of DocVQA (Mathew et al., 2021) marked a significant shift in focus from natural scene images to text-dense, layout-aware single-page document images, establishing a benchmark in the field. As advancements have continued, demands have recently emerged for models capable of addressing more complex, multi-page scenarios (Tito et al., 2023; Ding et al., 2023b, 2024a). These emerging requirements highlight the need for models to process multimodal inputs and navigate through extensive documents, reflecting user inquiries’ evolving complexity and naturalness in document-based question-answering systems. This section will briefly review the SoTAs in single-page document VQA models and introduce some recently proposed multi-page document understanding solutions (Tito et al., 2023; Blau et al., 2024).
3.3.1. Single-page VRD-QA
Similar to key information extraction from visually rich documents, single-page question answering (QA) often utilises classical pretrained language models (Devlin, 2018; Liu et al., 2019b) as baselines. These models conduct span-based question answering to extract sequences of text tokens. Additionally, general domain visual language pretrained models (Li et al., 2019; Tan and Bansal, 2019; Kim et al., 2021) are employed to identify the target document’s semantic entities (Ding et al., 2023b). Beyond these plain text or general domain vision language, numerous layout-aware models specifically pretrained in the document domain (see Section 3) have achieved state-of-the-art (SoTA) performance in single-page document tasks. This approach underscores the importance of integrating layout awareness in the preprocessing stages to enhance performance on these specific document-based tasks.
3.3.2. Multi-page VRD-QA
As demand increases for retrieving answers from multi-page documents (Tito et al., 2023; Ding et al., 2024a), most SoTA models (Xu et al., 2020a; Wang et al., 2022b; Hong et al., 2021), which are typically designed for single-page inputs, face a significant challenge due to their input length limitation of 512 tokens. To overcome this limitation, recent innovations have introduced solutions such as transformers capable of handling longer sequences and page-locating modules. These advancements are specifically designed to address the requirements of multipage document question answering (QA), enabling more efficient processing.
Hi-VT5 (Tito et al., 2023) proposes a multimodal hierarchical encoder-decoder architecture for multi-page generative question-answering. A T5-based multimodal transformer is applied to encode the single-page level information, including questions, OCR-extracted page content, image patches, and a set of page tokens. The question and OCR extracted token sequence is the following (Biten et al., 2022) to acquire the layout-aware initial textual representations. Document Image Trasnfoer (DIT) (Li et al., 2022b) is leveraged to acquire initial patch representations. Then, the concatenated question ($Q$) and page OCR tokens ($T$), image patch tokens ($I$), and randomly initialised page tokens ($P$) are fed into T5-based page encoder to enhance the multi-level and multimodal contextual learning contextually. The enhanced page token embeddings $P^{\prime}$ of each input document image are fed into a T5-based decoder to generate the predicted answer, as well auto-regressively, and the page located on the target answer must output the page number based on $P^{\prime}$. As T5 is not a layout-aware language model, masked language modelling is applied to predicted masked tokens by leveraging non-based visual and layout information to boost multimodal understanding. As T5 can accept input sequence lengths up to 20,480 tokens, it dramatically increases the standard VRDU pretrained models that apply Hi-VT5 to multi-page scenarios.
GRAM (Blau et al., 2024): proposes a framework to extend single-page models to tackle multi-page document VQA scenarios. Following the way of pretrained DocFormerv2 (Appalaraju et al., 2024), the framework encodes single-page inputs, including questions, OCR-extracted content, and visual features. For multi-page scenarios, a slim global encoder follows each single-page encoder layer, allowing the new learnable page tokens to interact with other pages contextually. This enables page tokens to capture page-level information through self-attention and enhance document-level understanding with sparse attention. As the global layer is a new stream applied to the pretrained model, an attention bias is applied following ALiBi (Press et al., 2021) to prevent the model from possibly disregarding the page tokens, enabling them to capture more fine-grained information from pretrained weights. During the decoding stage, unlike Hi-VT5, which only feeds page tokens to the decoder, GRAM uses all fine-trained information for the decoder. C-Former (Raffel et al., 2020) is applied to alleviate the high computation consumption, which could revise the information across all pages and distil only important details.
Kang et al. (Kang et al., 2024) introduce a multi-page document VQA framework to use a scoring self-attention mechanism to select and identify the related pages for generating the answer to the input question. The training processes include single-page Document VQA training and then training a self-attention scoring module based on the frozen trained encoder to feed the most relevant page information into the decoder for answer generating. Pix2Struct (Lee et al., 2022a) is used as the single-page model fine-tuned on the DocVQA (Mathew et al., 2021) dataset. The output from the Pix2Struct is fed into the self-attention scoring module to extract the first token for predicting a question-page matching score. The page with the highest matching score will be fed into the decoder for answer generation.
3.3.3. Summary
Document Visual Question Answering (Document VQA) is a relatively new field, pioneered by DocVQA (Mathew et al., 2021), which aims to generate or extract answers to natural language questions based on document images. This differs from key information extraction, which focuses on recognizing or extracting predefined key-value pairs. Document VQA requires a more comprehensive representation of the document and an understanding of correlations between the document and questions. Pretrained VRDU models (Xu et al., 2020b; Huang et al., 2022) demonstrate robust performance in single-page document understanding tasks. However, these models encounter challenges when applied to more typical and natural multi-page scenarios due to input length limitations. Recent solutions, such as (Tito et al., 2023; Blau et al., 2024), mainly address these challenges by identifying the page where a possible answer may be located and then applying SoTA single-page techniques to retrieve the answer. Despite these advancements, real-world applications often present more complex situations, such as long-term dependencies and cross-page entity relationships, which still need further exploration in the document VQA field.
4. Multi-Task VRD Understanding Models
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2 [Multi-task based models, for tree=fill=brown!45 [Fine-grained Pretrained Models (Encoder-only), for tree=fill=violet!40 [Layout-aware Pretrained Models, for tree=fill=violet!20 [LayoutLM (Xu et al., 2020a); BROS(Hong et al., 2021); LiLT (Wang et al., 2022b); XDoc (Chen et al., 2022); LayoutMask (Tu et al., 2023); StructuralLM (Xu et al., 2020a), for tree=fill=violet!10] ] [Visual-integrated Pretrained Models, for tree=fill=violet!20 [LayoutLMv2(Xu et al., 2020b); LayoutXLM(Xu et al., 2021); DocFormer(Appalaraju et al., 2021); LayoutLMv3(Huang et al., 2022), for tree=fill=violet!10] ] ] [Coarse and Joint-grained Pretrained Models (Encoder-only), for tree=fill=pink!40 [Coarse-grained Models, for tree=fill=pink!20 [SelfDoc(Li et al., 2021b); UniDoc(Gu et al., 2021)], for tree=fill=pink!10 ] [Joint-grained Models, for tree=fill=pink!20 [StrucText(Li et al., 2021c); Fast-StrucText(Zhai et al., 2023); MGDoc (Wang et al., 2022a); WUKONG-READER (Bai et al., 2022); GeoLayoutLM (Luo et al., 2023), for tree=fill=pink!10] ] ] [Encoder-Decoder Pretrained Frameworks, for tree=fill=red!40 [OCR-dependent Encoder-Decoder Frameworks, for tree=fill=red!20 [TiLT(Li et al., 2021b); UDOP(Tang et al., 2022); DocFormerv2(Appalaraju et al., 2024); ViTLP(Mao et al., 2024), for tree=fill=red!10] ] [OCR-independent Frameworks, for tree=fill=red!20 [Donut(Kim et al., 2022); Dessurt(Davis et al., 2022); ReRum (Cao et al., 2023a); StructTextV2 (Yu et al., 2022), for tree=fill=red!10] ] [LLM-based Frameworks, for tree=fill=red!20 [HRVDA(Liu et al., 2024b); LayoutLLM(Luo et al., 2024), for tree=fill=red!10] ] ] [Non-Pretrained Frameworks, for tree=fill=yellow!40 [CALM(Du et al., 2022); LayoutGCN(Shi et al., 2023),for tree=fill=yellow!20 ] ] ]
Figure 3. Multi-task visually rich document understanding frameworks.
Models designed for specific VRDU tasks often incorporate task-oriented techniques. Drawing inspiration from pretrained models in the vision (Dosovitskiy et al., 2020; Ren et al., 2015) and language (Devlin, 2018; Liu et al., 2019b) domains, enhancing document representations may significantly improve the performance of various downstream tasks. Consequently, a range of models have been developed to extract knowledge from extensive document collections. These models employ different pretrained tasks based on their architecture, the modalities they process, and the granularity of the information they extract from documents. Moreover, some models introduce specialised techniques to boost the effectiveness of pretrained model representations, achieving better performance without heavy pretraining. This section explores various document understanding frameworks focused on enhancing document representation for robustness and comprehensiveness in addressing multiple VRDU downstream tasks.
4.1. Fine-grained Pretrained Models
Inspired by BERT-style pretrained models, many researchers have proposed effective methods to integrate layout and visual information into models, aiming to enhance the comprehensiveness of textual token representations.
4.1.1. Layout-aware Pretrained Language Modeling
Understanding layout structure and spatial correlations between textual tokens can yield a more comprehensive representation of documents, going beyond what plain-text input offers. To this end, various methods have been proposed to encode layout features. These methods, combined with tailored pretraining tasks, enable models to capture layout-aware information better and effectively fuse textual and layout features.
LayoutLM (Xu et al., 2020a) is the first pretrained document understanding model by leveraging textual and layout information in the pretraining stage. BERT architecture is the backbone and 2-D positional embedding  [^3] Please refer to Section xx for more detailed information on 2-D positional encoding with textual information is used to pretrain on IIT-CDIP Test Collection 1.0. Two specific pretraining tasks are first introduced, named Masked Visual-Language Model (MVLM) and Multi-label Document Classification, to generate layout-aware textual representation and more comprehensive document representation, respectively. Like Masked Language Modeling adopted by most pretrained language models, MVLM randomly masks some input tokens but keeps the corresponding 2-D position embeddings to predict masked tokens to ensure the pretrained model is aware of the spatial relations between input tokens. MDC is a supervised pretraining task to predict the input document types (e.g. forms, exam papers, academic papers) that generate more comprehensive document-level representations. The fine-tuned LayoutLM could perform much better than textual-only frameworks on key information extraction (Jaume et al., 2019; Huang et al., 2019) and document classification (Harley et al., 2015).
BROS (Hong et al., 2021) aims to propose a pretrained VRDU model to represent the continuous property of 2D space by introducing a new 2-D positional encoding and a textual-spatial correlation aware attention score to replace the vanilla self-attention. Supposing $[x1,y1,x2,y2,x3,y3,x4,y4]$ is the bounding box coordinates of input token $t$, $p^{1}=[\mathcal{F}_{sin}(x1)\oplus\mathcal{F}_{sin}(y1)],p^{1}\rightarrow\mathbb{R}^{d_{pos}}$. The final positional representation of $t$ is $pos_{t}=W_{p1}p_{1}+W_{p2}p_{2}+W_{p3}p_{3}+W_{p4}p_{4}$, where all $W_{p}\in\mathbb{R}^{2d_{pos}\times d}$. The new 2-D positional encoding intends to provide a more natural way to encode the continuous 2-D coordinates. Additionally, instead of simply summing textual and positional representations, a novel attention score is introduced to consider both textual and spatial features and their correlations. To calculate the attention score $\alpha_{ij}$ between $t_{i}$ and $t_{j}$, they consider correlation intra and inter-correlation between textual and positional modalities.
(12) $\alpha_{ij}=(W^{q_{t_{i}}}T_{i})^{\top}(W^{q_{t_{j}}}T_{j})+(W^{q_{t_{i}}}T_{i}\circ W^{pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{pos_{t_{j}}}pos_{t_{j}})+(W^{\prime pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{\prime pos_{t_{j}}}pos_{t_{j}})$
where $(W^{q_{t_{i}}}T_{i})^{\top}(W^{q_{t_{j}}}T_{j})$ and $(W^{\prime pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{\prime pos_{t_{j}}}pos_{t_{j}})$ is the intra-modality attention scores of textual and positional modalities between two tokens. $(W^{q_{t_{i}}}T_{i}\circ W^{pos_{t_{i}}}pos_{t_{i}})^{\top}(W^{pos_{t_{j}}}pos_{t_{j}})$ is used to formulate the spatial dependency given the source semantic representation. Additionally, inspired by SpanBERT (Joshi et al., 2020), they use an area-masked language model to mask tokens by random-sized rectangle regions.
StructuralLM (Li et al., 2021a) is the first VRUD model using image patches (named ”cells” in the paper) to group input tokens for conducting various pretraining tasks. They use BERT as the backbone and generate multimodal representations $P$ of each patch, which will be used for two pretraining tasks, MVLM and Cell (Patch) Position Classification (CPC). The bounding box $(x0,y0,x1,y1)$ of an image patch is encoded by the 2-D positional encoding introduced by LayoutLM and the $n$ tokens inside $p$ are represented as $\{t_{1},t_{2},\dots,t_{n}\}$ which share the same 2-D positional encoding as for tokens belongs to one patch. A token $t_{i}$ could be represented as summing up token representation $T_{i}$, 2-D positional encoding $pos_{t_{i}}^{2d}$ and 1-D position embedding $pos_{i}$. Two pretraining tasks are used to understand the patch-level spatial correlations, including MVLM and CPC. MVLM is similar to LayoutLM but uses patch-level layout embeddings instead of token-level. Another novel cell position classification task is applied to predict the area index of randomly selected tokens from $N$ evenly split areas. Two pretraining tasks are performed simultaneously to capture the patch-level spatial dependencies between input tokens.
LiLT (Wang et al., 2022b) introduces a language-independent layout Transformer for mono/multi-lingual document understanding. The text and layout information are first separately encoded and jointly learned during pretraining. In the fine-tuning stage, two modality representations are concatenated to perform downstream tasks. The input token representations follow BERT, where $j$-th token is $T_{j}=T_{j}+pos_{t_{j}}+pos_{t_{j}}^{2d}$. The layout representations are slightly different from LayoutLM, normalising the bbox coordinates in the $[0,1000]$ and four linear layers encode the x-axis, y-axis, width and height features. The normalised bbox of a token $t$ is $[x0,y0,x1,y1,w,h]$ is encoded to get the final layout representation $L$:
(13) $L=W_{L}(W_{x}x0\oplus W_{y}x0\oplus W_{x}x1\oplus W_{y}y1\oplus W_{w}w\oplus W_{h}h)+pos_{L}$
where $W_{L}\in\mathbb{R}^{6d_{L}^{\prime}\times d_{L}}$ and $W_{x},W_{y},W_{w},W_{h}\in\mathbb{R}^{1\times d_{L}}$. The text and layout embedding are fed into two sub-models to generate high-level representations. A bi-directional attention complementation mechanism (BiACM) is proposed to augment the cross-modality interaction  [^4] We provide a detailed explanation of BiACM in Section X. Three self-supervised learning methods are proposed to enable the model to understand the multimodal document representation, including MVLM and Key Point Location, which is similar to the CPC proposed by StructureLM to predict the area index of masked tokens based on the layout features. Additionally, a text and layout alignment task is proposed to enhance further the cross-modality representation to predict whether each pair is aligned. LiLT could achieve promising performance on multi-lingual document understanding benchmarks (Wang et al., 2021a; Xu et al., 2021).
XDoc (Chen et al., 2022) propose a unified framework to deal with text inputs from multi-format inputs, including plain text, document and web text. Various encoding methods are proposed to tackle diverse text formats. For plain text token $t_{pln}$, the token representation follows BERT to get $T_{pln}$. For document text token $t_{doc}$, they adopted similar strategies (Xu et al., 2020a; Hong et al., 2021) to get $T_{doc}$ by summing up initial token representation $E_{t_{doc}}$, 1-D position embedding $pos_{t_{doc}}$ and 2-D position embedding $pos_{t_{doc}}^{2d}$. $pos_{t_{doc}}^{2d}$ uses different bbox formats $[x_{0},x_{1},y_{0},y_{1},w,h]$ following a Linear-ReLu-Linear adaptor to make the document format text embedding more representative. To encode web text inputs, they follow the way introduced by MarkupLM to encode the source file tags and subscript information to get the XPath embedding. The same structured adaptor is leveraged for better pretraining. All three text formats are pretrained by MLM and fine-tuned on plain text (Rajpurkar et al., 2016a; Wang et al., 2018), document (Jaume et al., 2019; Mathew et al., 2021), and web text (Chen et al., 2021) benchmark datasets, respectively.
LayoutMask (Tu et al., 2023) aims to improve the text-layout interactions by using local 1D positional encoding and new pretraining tasks. OCR-dependent document understanding frameworks are suffered from improper reading orders. It uses LayoutLMv2 as the backbone but removes the visual module, and the local 1D positional encoding is adopted to replace the global 1D positional encoding, which only encodes the token ordered within each segment detected by OCR tools and always restarts with 1 from each individual segment. Moreover, the 2D positional encoding is also using segment bbox instead of individual words. The first task pertains to traditional masked language modelling, but two different masking strategies are used. Firstly, they set the mask at the word level instead of masking the token itself to enable the model to capture more contextual information to predict the masked word. Additionally, to boost cross-segment understanding, the probability of masking the first and last word of a segment is higher than that of others. Another Masked Position Modelling is proposed to ask the model to predict the masked word-level 2D positions to promote the layout information representation.
4.1.2. Visual Integrated Models
Integrating visual cues alongside text and layout information during the pretraining stage can significantly enhance model capabilities, capturing more comprehensive document insights than text and layout alone. Various frameworks and pretraining tasks focusing solely on text and layout have been expanded to include visual-text matching tasks to strengthen cross-modality alignment. This integration enables models to interpret better the complex interplay of visual, textual, and layout features within documents. Existing models that utilise these comprehensive inputs can be categorised based on their approach to visual feature acquisition: feature map-based models, which generate comprehensive visual representations, and patch pixel-based models, which focus on granular visual details. This classification helps understand how different models leverage visual information to enhance document understanding.
LayoutLMv2 (Xu et al., 2020b) is the first pretrained model integrating text, layout and visual aspects during the pretraining stage. A single framework multimodal transformer is applied to get the visual, text and layout features simultaneously, where each input textual and visual tokens are assigned a segment ID to distinguish the modality or semantic type. The textual and layout representations generally follow LayoutLM, except the linear projected segment ID, $seg_{t}$, embeddings are addicted to original textual embeddings. The input document image is firstly fed into a trainable visual encoder (based on ResNeXt-FPN), of which output is evenly split into $m$ flattened visual tokens, $v_{token}$ following a linear layer to project into the same dimension as textual embedding. The $j$-th visual token embedding is $V_{j}=W_{v_{j}}+pos_{j}+W_{seg_{v}}$. Moreover, to capture the relative position information of inter/intra-modality features, spatial-aware self-attention is introduced by adding bias terms of 1-D ($b^{1D}$) and 2-D ($b^{2D}$) relative position. The 1-D relative positional bias between input visual or textual tokens $t_{i}$ and $t_{j}$ is $b^{1D}=W_{b_{1D}}(j-i)$ and $b^{2D}=b^{2D}_{x}+b^{2D}_{y}$, where $b^{2D}_{x}=W_{b^{2D}_{x}}(x0_{i}-x0_{j})$, $b^{2D}_{y}=W_{b^{2D}_{y}}(y0_{i}-y0_{j})$. LayoutLMv2 is pretrained on MVLM and the other two new proposed tasks, Text-Image Alignment (TIA) and Text-Image Matching (TIM)  [^5] Please refer to Section to check more detailed information about those two pretraining tasks. to capture more cross-modality alignment.
LayoutXLM (Xu et al., 2021) extends the LayoutLMv2 architecture to a multilingual setup where the MVLM is extended to Multilingual Mased Visual-Language Modeling pretrained on 22 million self-collected digital-born PDF files and 8 million scanned English documents from IIT-CDIP processed by off-the-shelf PDF parsers.
DocFormer (Appalaraju et al., 2021) is a multimodal transformer encoder-based architecture which also uses a trainable CNN-backbone (ResNet50) to extract the visual cues of input document image to get visual representations $\mathbb{V}\in\mathbb{R}^{(d\times N)}$, where $d=768$ is the transformer hidden state and $N=512$ is the number of visual tokens. The initial textual representations $\mathbb{T}$ are acquired by feeding the OCR extracted text with bbox into LayoutLM where $\mathbb{T}\in\mathbb{R}^{(d\times N)}$. 2D positional encoding is also used by using $W^{x}_{v},W^{y}_{v}$ and $W^{x}_{t},W^{y}_{t}$ to encode $x,y$ axis, as well as, visual and textual aspects to acquire $pos^{2d}_{v}$ and $pos^{2d}_{t}$. More positional features are adopted such as width, height and relative distance between neighbours. Moreover, certain inductive biases are applied to get a new self-attention score $\alpha$ for paying more attention to local features. Supposing $\alpha_{ij}^{v}$ is attention score between visual feature $V_{i}$ and $V_{j}$, we could have $\alpha_{ij}^{v}=(W^{K}_{v}V_{j})^{\top}(W^{Q}_{v}V_{i})+(pos_{ij})^{\top}W^{Q}_{v}V_{i}+(pos_{ij})^{\top}W^{K}_{v}V_{j}+(W^{Q}_{s}pos_{ij}^{2d})^{\top}(W^{K}_{s}pos_{ij}^{2d})$, where $pos_{ij}$ is the 1D relative positional encoding. The same procedures are applied to textual features as well, and two modalities share the same $W^{Q}_{s}$ and $W^{K}_{s}$ to help the model correlate features across modalities. The multi-modal token representations $\mathbb{M}$ of $l$ encoder layer is $\mathbb{M}_{l}=\mathbb{T}_{l}+\mathbb{V}_{l}$.
LayoutLMv3 (Huang et al., 2022) is the first pretrained VRDU model to encode visual features without using heavy CNN-backbones. LayoutLMv3 uses the exact same way as LayoutLMv2 to encode the textual and layout information except replacing the word-level bbox to segment-level for 2-D positional embedding. For visual feature representation, they follow ViT and ViLT to linearly project the evenly split image patches with learnable 1-D positional encoding to the transformer encoder. For the pretraining setting, they use masked language modelling to mask 30% tokens drawn from a Poisson distribution. To augment the visual representation by contextually learning with multimodal features, a Masked Image Modelling (MIM) (Bao et al., 2021) task is used to mask 40% image tokens with clockwise masking randomly. Each image token will be converted into discrete tokens following (Ramesh et al., 2021), and a cross-entropy loss is adopted to reconstruct the masked image tokens. Moreover, to explicitly learn the correlation between text and image modalities, a Word-Patch Alignment (WPA) pretraining objective is applied. WPA is a binary classification task to predict whether the unmasked token-patch pair is ”aligned” or ”unaligned”.
4.2. Coarse and Joint-grained Pretrained Models
Fine-grained models achieve state-of-the-art performance on many downstream tasks but face challenges with input length limitations and capturing document image layout and logical arrangement. To address these issues, coarse-grained or joint-grained frameworks have been introduced. To mitigate these limitations, these frameworks leverage multimodal information from document semantic entities such as paragraphs, tables, and textlines.
4.2.1. Coarse-grained Frameworks
SelfDoc (Li et al., 2021b) is the first pretrained VRDU model to leverage coarse-grained document elements for various document understanding tasks. Unlike fine-grained models that rely on OCR-extracted text sequences, SelfDoc uses Faster-RCNN to extract Regions of Interest (RoIs) from document entities. This approach reduces the input sequence length for text-dense and long documents with improved time and space complexities. To acquire the initial multimodal representations, Faster-RCNN extracts visual embeddings from these RoIs, while Sentence-BERT (Reimers and Gurevych, 2019) provides textual embeddings based on the OCR-extracted text of each entity. The textual and visual representations are fed into two single-modality BERT-like encoders to learn the intra-modality correlations between entities contextually. A cross-modality encoder is introduced to enhance inter-modality learning, mirroring the structure of single-modality encoders but incorporating cross-attention layers. It randomly masks entities in document images within either language or vision branches during pretraining. Furthermore, a modality-adaptive attention mechanism dynamically adjusts weights for visual and textual embeddings based on the input and task, creating robust entity representations for fine-tuning.
UniDoc (Gu et al., 2021) is an entity-level document understanding model which contains a trainable image encoder with RoI-Align (He et al., 2017) to extract entity visual representation and novel cross-attention mechanisms to fuse multimodal information. The initial sentence and textual embeddings are acquired by RoI features and averaged word embeddings summing up with the linear projected bbox coordinates. To enable the trainable visual representation to predict meaningful visual cues effectively, product quantization (Jegou et al., 2010) is applied to discretise a RoI feature into a finite set of visual representations and mapped to a new embedding. Additionally, a novel Gated Cross-Attention is introduced to improve the interactive learning between various modalities. After a multi-head cross-attention to get the enhanced visual and textual representations, a gating mechanism is applied to dynamically weight the visual ($V$) and textual features ($T$) by feeding concatenated $[V:T]$ to a non-linear network to acquire a modality-aware attention bias $\beta_{v},\beta_{t}$ for visual and textual respectively. There are three pretraining tasks are conducted simultaneously to enhance multimodal feature representatives. Masked Sentence Modelling and Visual Contrastive Learning are required to predict the masked sentence representation and quantize visual representation, respectively. Another Vision-Language Alignment introduced by LayoutLMv2 (Xu et al., 2020b) to enforce contextual learning between multi-modalities. But instead of using split image regions, UniDoc aligns the image and text to belonging to the same entity.
4.2.2. Joint-grained Frameworks
StrucText (Li et al., 2021c) is a multimodal pretrained VRDU model using fine-grained textual and coarse-grained vision information to capture richer geometric and semantic information from different levels and modalities. The layout embedding of each text token and segment is encoded by $L=W_{l}[x0,y0,x1,y1,w,h],W_{1}\in\mathbb{R}^{6\times N}$. They follow token-level models to encode the sequence of tokens and use pretrained ResNet50 with FPN to extract the entity visual features. A segment-ID embedding is allocated to each token’s textual and entity visual representations to boost the alignment learning. Three self-supervised tasks are introduced to enhance inter-modality learning: MVLM, Segment length Prediction (SLP), and Paired Box Direction (PBD). As a new self-supervised learning task, SLP asks the model to predict the entity’s length to leverage the entity’s visual embeddings and textual information from the same segment ID. Another self-supervised learning task aims to learn more comprehensive pair-wise spatial correlations between entities by clarifying their spatial correlation.
Fast-StructText (Zhai et al., 2023) is built on StrucText to improve model efficiency and enhance the feature expressiveness by introducing an hourglass transformer and Symmetry Cross-Attention mechanisms (SCA). Similar feature encoding methods to StructText are adopted, but the bbox formats of layout encoding are $[x0,y0,x1,y1]$. To encode the module and progressively reduce the redundancy tokens, an hourglass transformer is proposed consisting of several Merging blocks and Extension blocks to down-sampling and up-sampling the number of input tokens. Merging blocks suggest merging neighbouring $k$ tokens with the weighted 1D average pooling to shorten the sequence length. The extension block is required to transform the shortened sequence back to the entire length by simply applying a repeat up-sampling method. In addition, to enhance the interaction between textual and vision modalities, SCA consists of two dual cross-attention to handle text and visual features. Different self-supervised tasks are conducted simultaneously, including MVLM (Xu et al., 2020b; Li et al., 2021c), Graph-based Token Relation (GTR), Sentence Order Prediction (SOP) and Text-Image Alignment (Huang et al., 2022). For the two new proposed tasks, GTR is a task similar to Paired-Box Direction to predict the positional relations between paired entity visual features and SOP is used to predict whether two sentence pairs are normal-order adjacent to learning semantic knowledge.
MGDoc (Wang et al., 2022a) is the first multimodal multi-granular pretrained framework which introduces multi-granular attention and cross-modal attention to increase the inter-grained learning and better cross-modality fusion. For acquiring initial features of different modalities, pretrained language models (Reimers and Gurevych, 2019) and vision encoders (He et al., 2016) are used to encode different-grained inputs from word to whole page. The encoded modalities are associated with positional features (Zhai et al., 2023) and modality-type embeddings (Li et al., 2021c) to acquire the final input representation of various modalities. To encode the hierarchical relation between multi-grained features, including pages, entities and words, two attention biases are added to the original self-attention weights. A hierarchical bias is a binary value of 0 or 1 to examine the inside or outside relation between two inputs, while the relation bias is the relative positive between two bboxes. Apart from multi-grained learning, cross-attention is applied to fuse cross-modality information better. Three tasks are proposed during pretraining, including Mask Text Modeling (MTM), Mask Vision Modeling (MVM) and Multi-Granularity Modeling. MTM and MVM randomly mask the multi-grained textual and visual representations to predict the masked features under the mean absolute error. To boost the interactive understanding of multi-granularity, a token-entity linking prediction is proposed by computing the dot-product between token and entity features.
WUKONG-READER (Bai et al., 2022) uses fine-grained level inputs but leverages coarse-grained level self-supervised tasks to enhance the fine-grained level information. The model inputs contain a document image and OCR extracted sequence of tokens with their bboxes. A Mask-RCNN is used as a visual backbone to acquire several visual tokens and extract visual features of textlines from a RoIHead layer. For textual information encoding, the first is layers of RoBERTa(Wei et al., 2020) are applied to extract the token representation and sum up with additional features following (Xu et al., 2020b). The concatenated visual and textual features are fed into a multimodal modal encoder based on the rest six layers of RoBERTa. Different pretraining tasks are proposed, including MLM, Textline-Region Contrastive Learning (TRC), Masked Region Modeling (MRM) and Textline Grid Matching (TCM). TRC aims to enhance cross-modality interactive learning at the entity (textline) level by following a text-vision-aligned contrastive learning approach (Yao et al., 2021). To enhance the visual representation, MRM is applied to mask 15% of textlines randomly and to predict the masked visual embeddings following (Li et al., 2021b). A text-grid alignment task is introduced to enhance layout understanding by dividing the document image into N-grids and predicting the tokens in 15% selected textlines belonging to which grid. The scaling parameters are applied to each loss of pretraining tasks.
GeoLayoutLM (Luo et al., 2023) is a sophisticated multimodal framework that distinctively incorporates geometric information through specialised pretraining tasks and the development of innovative relation heads. Inspired by the dual-stream structure of METER and SelfDoc (Li et al., 2021b), GeoLayoutLM features separate vision and text-layout modules coupled with interactive co-attention layers that enhance the integration of visual and textual data. The model introduces two advanced relation heads—the Coarse Relation Prediction (CRP) head and the Relation Feature Enhancement (RFE) head—which refine relation feature representation crucial for both pretraining and fine-tuning phases. The pretraining regimen includes tasks designed to understand geometric relationships, such as GeoPair, GeoMPair, and GeoTriplet, aiding the model in grasping the complex dynamics of document layouts. During fine-tuning, the model utilises pretrained parameters to optimise both semantic entity recognition and relation extraction tasks, employing a novel inference technique that enhances relation pair selection accuracy by focusing on the most probable relationships and minimizing variance among potential options.
4.3. Encoder-Decoder Pretrained Frameworks
In addition to the above encoder-only frameworks, researchers have proposed encoder-decoder pretrained models that often approach tasks like Key Information Extraction (KIE) or Visual Question Answering (VQA) in a generative style. Addressing limitations of OCR-dependent frameworks, such as accumulated OCR errors and incorrect reading orders, OCR-independent models have been introduced for end-to-end VRDU.
4.3.1. OCR-dependent Encoder-Decoder Frameworks
TILT (Powalski et al., 2021) is a T5-based transformer encoder-decoder architecture enhanced with a relative spatial bias in the self-attention mechanism to acquire fine-grained token representations. It incorporates encoding methods that apply relative sequence input bias and capture horizontal and vertical distance biases in the attention scores. A U-Net-based framework is applied to extract the fixed-size feature maps fed into the encoder together. They follow T5 pretraining strategies on RVL-CIDP dataset (Harley et al., 2015) but use a salient span masking scheme adopted by (Raffel et al., 2020; Guu et al., 2020). As the first encoder-decoder framework, TILT requires off-the-shelf OCR tools to acquire the textual token sequence. To conduct VRD content understanding end-to-end, some OCR-free frameworks are proposed to solve the limitations of OCR-dependent models.
UDOP (Tang et al., 2022) proposes an encoder-decoder pre-trained document understanding framework that leverages a ViT-based model (Dosovitskiy et al., 2020), following the principles of LayoutLMv3 (Huang et al., 2022), to process multimodal information. To enhance the comprehensiveness of textual token embeddings, the framework sums the text token embeddings with token-aligned image patch embeddings when the centre of the bounding box falls within the image patch. Additionally, the positional biases introduced by TILT are added, but ID positional encoding is not used. The decoding stage is designed to generate all vision, text and layout modalities consisting of a bi-directional Transformer text-layout decoder and an MAE vision decoder. Two decoders are cross-attended with each other. The pertaining tasks contain self-supervised on unlabeled documents to learn robust document representation and supervised pertaining tasks for fine-grained model supervision. Masked text layout and image reconstruction are used to predict the masked information using unmasked multimodal information. A layout moulding task is applied to predict the positions of groups of text tokens, and a visual text recognition task is proposed to identify text at the given location in the image. Supervised tasks utilise a training set of publicly available benchmark datasets of different downstream tasks, including document classification (Harley et al., 2015), KIE (Stanisławek et al., 2021), VQA (Tanaka et al., 2021; Mathew et al., 2021) and layout analysing (Zhong et al., 2019).
DocFormerv2 (Appalaraju et al., 2024) is an encoder-decoder transformer architecture that uses multimodal (visual, textual and positional) to enhance the multimodal understanding and layout-aware language decoder to predict the predictions. The patched image pixels are fed into the convolutional and linear layers to get down-sampled patch embedding. The textual embeddings are acquired by linear projected token one-hot encoding. Both visual and textual embeddings are summed with the 2D-positional encoding (Xu et al., 2020a) of Patches and linear project bbox embedding ($[x0,y0,x1,y1]$) (Wang et al., 2022b), respectively. Two encoder-based Token to Line (T2L), Token to Grid (T2G), and one decoder-based self-supervised learning task, MLM, are proposed to enable multimodal feature interactive learning. T2L aims to improve the relative position understanding between tokens by predicting the number of textlines between two randomly selected tokens. For improving the layout and structure, understanding needs to split the image into $m\times n$ grids to predict the located grid number of each OCR token. For the decoder MLM, the spatial feature of each masked text token is masked as well, and the other setup follows T5 (Raffel et al., 2020).
ViTLP (Mao et al., 2024) introduces an encoder-decoder architecture for OCR and document understanding using a ViT-based vision encoder to obtain image patch representations, which are fed into a decoder to generate text and layout sequences autoregressively. A special ”[LOC]” token, encoding bounding box coordinates $[x0,y0,x1,y1]$, reduces the layout token sequence length. To control generation flow, ”[BOS]” and ”[CONT]” tokens are added, representing input sequences of two tokens, t1,t2, as ”[BOS], t1, [LOC], t2, [LOC]”. The decoder has hierarchical heads: the text head uses all tokens to generate the next text token, while the layout head uses ”[LOC]” tokens to predict bounding box coordinates. The ”[CONT]” token handles sequences of arbitrary length by continuing generation until ”[END]”, based on the prefix token ratio.
4.3.2. OCR-free Pretrained Frameworks
[…]
ReRum (Cao et al., 2023a) introduce an end-to-end architecture in which the decoding process focuses on local interest and visual cues. Like other OCR-free frameworks, a Swin-Transformer extracts image patch representations. The extracted visual features are fed into a transformer-based Query-Decoder to acquire the vision-enhanced query representation. The enhanced query representation with visual features is fed into a Text Decoder to generate the predicted text tokens auto-regressively. Notably, a Content-aware Token Merge technique is proposed to dynamically focus on more relevant foreground parts of visual features, which selects top-K visual tokens based on the averaged correlation scores between query and visual token representations. Additionally, the unselected visual tokens (called Background Area) contain rich global features that could be used to enhance the Top-K foreground visual tokens through basic attention mechanisms. Three pretraining tasks are applied: query to segmentation (Q2S), text to segmentation (T2S), and segmentation to text (S2T). Q2S follows DETR setup for a token generation but alters to an instance segmentation task to predict $N$ mask for the target text area to improve text detection ability. T2S acquire the Text Decoder output to conduct another instance segmentation task after feeding the image text snippet into the Text Decoder to boost the layout-aware textual information understanding. The last S2T is to use the outputs from the Text Decoder to generate the token like an OCR task auto-regressively.
StrucTextV2 (Yu et al., 2022) is an end-to-end structure that uses image-only input to conduct several downstream tasks. It contains a CNN-based visual extractor with FPN strategies (Lin et al., 2017) and follows ViT (Dosovitskiy et al., 2020) to get linear projected flattened patch-level representations. The patch token embeddings serve as the input to the Transformer encoder to enhance the contextually semantic representations. Then, the lightweight fusion network is applied to generate the final representations and fed into two branches during pretraining: made language Modeling (MLM) and Masked Image Modeling (MIM). Instead of using text inputs when MLM is used by other models (Devlin, 2018), a portion of the text regions are masked with RGB values $[255,255,255]$ randomly with a 2-layer MLP decoder to predict the masked token. MIM masks the rectangular text regions and predicts the RGB values of the missing pixels to improve the document representations. Except for the global average pooled FPN fused visual representations, the MLM-generated hidden state of each text region is concatenated and fed into a Fully Convolutional New York to get the regressed masked missing pixel values.
4.3.3. LLM-based Frameworks
HRVDA (Liu et al., 2024b) aims to propose a MLLM accepting high-resolution image inputs to conduct fine-grained information extraction from VRDs. A swin-transformer (Liu et al., 2021) is used to encode document images into image patch tokens. A pluggable content detector then identifies visual tokens that contain relevant document content information. Following this, a content filtering mechanism performs token pruning to remove irrelevant tokens. The remaining encoded visual tokens are processed through an MLP to ensure consistency with the LLM embedding space dimensions. These pruned tokens are then fused with instruction features, allowing further filtering of tokens irrelevant to the instructions. The final streamlined set of visual tokens and instructions is fed into the LLM, which generates the corresponding responses.
LayoutLLM (Luo et al., 2024) introduces an LLM/MLLM-based approach integrated with a pretrained document understanding model to address the challenges of applying LLMs to zero-shot document understanding tasks. The input document’s visual, textual, and layout information and any question text are encoded by a pretrained LayoutLMv3 (Huang et al., 2022) encoder and projected into the same embedding space as the adopted LLM, Vicuna-7B-v1.5 (Zheng et al., 2024). The method incorporates layout-aware pretraining tasks at three levels: document-level (e.g., document summarization), region-level (e.g., layout analysis), and segment-level (e.g., MVLM). These tasks enable the model to achieve comprehensive document understanding. Additionally, a novel module called LayoutCoT is designed to help LayoutLLM focus on question-relevant regions and generate accurate answers through intermediate steps. GPT-3.5-turbo (OpenAI, 2023) is used to prepare the dataset for document summarization training and to construct LayoutCoT training data.
4.4. Non-Pretrained Frameworks
CALM (Du et al., 2022) introduces a common-sense augment document understanding framework to understand the query and extrapolate answers not contained in the context of the input document image. They follow LayoutLMv2 (Xu et al., 2020b) to encode input document multimodal representations. The textual token embeddings are fed into a Document Purifier component to merge the tokens $\{t_{1},\dots,t_{n}\}$ belonging to one entity type $N$ to one Upper Layer token $\hat{c}$ by applying average pooling of $\hat{c}=AvePool(t_{1},\dots,t_{n})$. Each Upper Layer token is concatenated with the commence augmented on ConceptNet NumberBatch (Speer et al., 2017) entity word vector $c^{\prime}$ to get the final entity representation $c=concat(\hat{c},c^{\prime})$. A similar Question-Purifier is applied to use common-sense knowledge to enhance the question representation. Then, with the assistance of ConceptNet, relevant common-sense knowledge is recalled based on the common-sense representation of both documents and queries. By considering the predicted question-answer relationship, a final self-attentive graph convolutional network following (Zhu et al., 2021b) is proposed to address document reasoning tasks more effectively.
LayoutGCN (Shi et al., 2023) proposes a lightweight and effective model which contains a fully connected graph where text blocks are nodes and edges connect every two blocks. The model architecture includes a TextCNN-based (Kim, 2014) encoder to encode N-gram textual embeddings, a linear trainable layout encoder to project the normalised bbox coordinates into hyperspace following other layout-aware models, and a visual encoder (CSP-Darknet (Wang et al., 2020a) for document image features). These features are integrated using a Graph Convolution Network (GCN) to capture relationships between nodes. The final node representation combines text, layout, and visual information, benefiting various VRDU tasks.
XYLayoutLM (Gu et al., 2022) introduces an Augmented XY Cut module to correct the improper reading order generated by OCR engines and a Dialted Conditional Position Encoding module to handle the variable lengths of input text and image tokens. LayoutXLM (Xu et al., 2021)  [^6] Please refer to Section 4.1.2 to see a detailed description of the LayoutXLM model. is adopted as a basic framework to process multimodal inputs. The Augmented XY Cut module enhances traditional XY Cut (Nagy and Seth, 1984) by incorporating thresholds ($\lambda_{x},\lambda_{y}$) and a shift factor ($\theta$) to adjust box positions on the x and y axis based on randomly generated values. It recursively divides token boxes using valleys in horizontal and vertical projection profiles, forming clusters in descending order. Each cluster’s reading order is determined recursively, prioritizing divisions until no significant valleys remain, ensuring a proper reading order is derived from an XY Tree structure. DCPE (Dilated Conditional Position Encoding) addresses limitations of CPE (Chu et al., 2022) in multimodal tasks by separately processing textual and visual features. It employs 1D convolutions for textual tokens to extract 1D local layouts, accommodating their inherent 1D relationships. Additionally, DCPE utilizes dilated convolutions (Yu and Koltun, 2015) to capture long-range dependencies effectively without increasing model complexity, thereby enhancing performance in multimodal document understanding tasks.
4.5. Summary of Multi-Task Frameworks
Various models are proposed to enhance document representations for VRDU tasks by leveraging pretrained language models to enrich text token sequences with layout information through positional encoding (Xu et al., 2020a), attention mechanisms (Wang et al., 2022b), and layout-aware tasks (Tu et al., 2023). However, VRDs contain rich visual details like font, texture, and colour and visually complex entities such as tables, charts, and photos. Many models (Xu et al., 2020b; Huang et al., 2022; Appalaraju et al., 2021) integrate visual cues to enhance fine-grained document features, but their quadratic time and space complexity pose challenges for handling long sequences in multi-page document understanding (Ding et al., 2024a). Fine-grained models excel but struggle with capturing layout and structural details from document images. Coarse-grained frameworks (Li et al., 2021b; Gu et al., 2021) mitigate fine-grained limitations by leveraging entity-level multimodal information, yet compressing diverse entity aspects into a single dense vector risks losing information (Ding et al., 2024b). Joint-grained frameworks (Li et al., 2021c; Zhai et al., 2023; Wang et al., 2022a; Bai et al., 2022; Luo et al., 2023) integrate multi-grained information to produce comprehensive representations. Non-pretrained models leverage external knowledge (Du et al., 2022) or lightweight networks (Shi et al., 2023) to rival large-scale pretrained frameworks’ performance. Most document understanding models (Xu et al., 2020a; Huang et al., 2022; Wang et al., 2022b; Li et al., 2021b) rely on off-the-shelf OCR tools for text extraction, which can be susceptible to OCR quality issues and incorrect reading orders. OCR-free frameworks directly process document images to mitigate these limitations; However, these frameworks may exhibit sub-optimal performance compared to methods using established OCR tools with additional resource consumption.
5. Visually Rich Document Content Understanding Datasets
5.1. Key Information Extraction and Entity Linking
Name Conf./J. Year Domain # Docs # Images # Keys MP. Language Metrics Format FUNSD ICDAR-w 2019 Multi-source N/A 199 4 N English F1 P.& H. SROIE ICDAR-c 2019 Scanned Receipts N/A 973 4 N English F1* P. CORD Neurips-w 2019 Scanned Receipts N/A 1,000 54 N English F1 P. Payment-Invoice ACL 2020 Invoice Form N/A 14,237+595 7 N English F1 D. Payment-Receipts ACL 2020 Scanned Receipts N/A 478 2 N English F1 P. Kleister-NDA ICDAR 2021 Private Agreements 540 3,229 4 Y English F1 D. Kleister-Charity ICDAR 2021 AFR 2,778 61,643 8 Y English F1 D.& P. EPHOIE AAAI 2021 Exam Paper N/A 1,494 10 N Chinese F1 P.& H. XFUND ACL 2022 Synthetic Forms N/A 1,393 4 N Multilingual F1 D.& P.& H. Form-NLU SIGIR 2023 Financial Form N/A 857 12 N English F1 D.& P.& H. VRDU-Regist. Form KDD 2023 Registration Form N/A 1,915 6 N English F1 D. VRDU-Ad-buy Form KDD 2023 Political Invoice Form N/A 641 9+1(5) N English F1 D.&P. DocILE ICDAR 2023 Invoice Form 6,680 106,680 55 Y English AP, CLEval D.&P.
Table 1. Summary of visually rich document key information extraction datasets.
5.1.1. Scanned Receipt Datasets
SROIE (Huang et al., 2019) is a widely used dataset for text localization, OCR, and key information extraction from scanned receipts, introduced in the ICDAR 2019 Challenge on ”Scanned Receipts OCR and Key Information Extraction” The Key Information Extraction (KIE) task focuses on four key types: Address, Date, Company, and Total, with corresponding values provided in the annotation file for each receipt. The F1 score for this task is calculated based on Mean Average Precision (MAP) and recall. Note that entity-level annotations are not provided in the official dataset, requiring the use of external tools to obtain the information for coarse-grained or joint-grained models.
Payment-Receipts (Majumder et al., 2020b) is a subset of SROIE, created by sampling up to 5 documents from each template in the original SROIE dataset. The template of each receipt is decided by the Company annotation. The target schema focuses on extracting only Date and Total. This subset is used to evaluate the model’s ability to handle unseen templates.
CORD (Park et al., 2019) is a widely used dataset for post-OCR receipt understanding, featuring two-level labels annotated by crowdsourcing workers. It includes eight superclasses, such as Store, Payment, Menu, Subtotal, and Total, each with several subclasses. For example, Store contains subclasses like Name, Address, and Telephone. CORD provides both textline-level and word-level annotations for both fine-grained and coarse-grained frameworks, with some sensitive information blurred. All models are evaluated on the released first 1,000 samples.
5.1.2. Form-style Datasets
FUNSD (Jaume et al., 2019) is derived from the RVL-CDIP dataset (Harley et al., 2015) by manually selecting 199 readable and diverse template form images. The dataset is annotated using the GuiZero library to provide both entity and word-level annotations, including manual text recognition. Semantic links indicate relationships between entities, such as Question-Answer or Header-Question pairs. Consequently, FUNSD supports key information extraction, OCR, and entity linking tasks.
XFUND (Xu et al., 2021) is the first multilingual dataset following the FUNSD format. It collects form templates in seven languages (Chinese, Japanese, Spanish, French, Italian, German, and Portuguese) from the internet. Human annotators fill these templates with synthetic information by typing or handwriting, ensuring each template is used only once. The filled forms are then scanned into document images, processed with OCR, and annotated with key-value pairs. Each language has 199 annotated forms, supporting multilingual key information extraction and entity linking tasks.
Payment-Invoice (Majumder et al., 2020b) contains two corpora of invoices from different sources. The first corpus, Invoice 1, includes 14,273 invoices from various vendors with different template styles, used for training and validation. The second corpus, Invoice 2, comprises 595 documents with distinct templates not found in Invoice 1, serving as the test set. Human annotators extract six required keys from each single-page invoice, such as Invoice Date, Total Amount, and Tax Amount. This dataset is suitable for evaluating generative-style models and fine-grained sequence labeling models. For coarse-grained models, additional text line or entity-level information can be extracted using off-the-shelf tools.
VRDU-Registration Form (Wang et al., 2023b) is a dataset of registration forms about foreign agents registering with the US government collected from the Federal Communications Commission. Commercial OCR tools extract the text content of the forms. Annotators draw bounding boxes around six unrepeated entities (each entity appears only once per document) per document: File Date, Foreign Principal Name, Registrant Name, Registration ID, Signer Name, and Signer Title. The dataset provides entity-level annotations, which can be easily preprocessed to acquire token-level annotations, supporting any granularity of Key Information Extraction (KIE) models.
VRDU-Ad-buy Form (Wang et al., 2023b) consists of 641 invoices or receipts signed between TV stations and campaign groups for political advertisements. It follows the same annotation procedure as the VRDU-Registration Form but involves a more complex schema. This includes nine unique entities (e.g., Advertiser, Agency, Contract ID), four repeated entities (e.g., Item Description, Sub Prices), and hierarchical entities (e.g., Line Item). Repeated entities may contain different values within a single document, while hierarchical entities comprise several repeated entities as components.
Form-NLU (Ding et al., 2023a) is a visual-linguistics dataset designed to support researchers in interpreting specific designer intentions amidst various types of noise from different form carriers, including digital, printed, and handwritten forms. Fine-grained key-value pairs, such as Company Name, Previous Notice Date, and Previous Shares, are manually annotated. The training and validation set comprises 535 digital-born forms, with 76 reserved for validation. Additionally, three test sets are provided, containing 146 digital, 50 printed, and 50 handwritten form images, respectively. Form-NLU can be used to evaluate form layout analysis and Key Information Extraction (KIE) models of any granularity. With proper processing, it can also be used to evaluate entity linking frameworks, thanks to the well-annotated key-value pairs.
5.1.3. Multi-page Datasets
Kleister-NDA (Stanisławek et al., 2021) is a dataset collected from the Electronic Data Gathering, Analysis, and Retrieval System (EDGAR) focusing on Non-disclosure Agreements (NDAs). During preprocessing, the collected 540 HTML files are converted into digital multi-page PDF files (totalling 3,229 pages) using the Puppeteer library. Four key items, Effective Date, Party, Jurisdiction, and Term, are manually annotated by three annotators to extract the corresponding values. The NDA dataset is widely used by fine-grained level models but may require additional processing for frameworks with limited sequence length due to the multi-page inputs.
Kleister-Charity (Stanisławek et al., 2021) contains 2,778 annual financial reports from the Charity Commission, which lack strict formatting rules. The Charity Commission website provides eight key pieces of information, such as Postcode, Charity Name, and Report Date. Annotators manually correct minor errors to ensure accuracy. Compared to Kleister-NDA, the Charity dataset has longer document inputs, totalling 61,643 pages, requiring models to handle long sequence outputs. Both Charity and NDA datasets provide only key-value pair annotations, making them suitable for the generation and fine-grained sequence labelling tasks but requiring additional processing to acquire entity-level annotations.
DocILE (Šimsa et al., 2023) comprises three subsets: an annotated set of 6,680 real business documents, an unlabeled set of 932,000 real business documents for unsupervised pretraining, and a synthetic set of 100,000 documents generated with full task labels. Documents come from public sources like the UCSF Industry Documents Library and Public Inspection Files, with annotations for Key Information Localization and Extraction and Line Item Recognition. Synthetic documents were created using annotated templates and a rule-based synthesizer. DocILE provides entity-level annotations that can be easily post-processed to acquire token-level annotations.
5.2. Visual Question Answering
Name Conf./J. Year Domain Lang. # Docs # Images # Questions Answer Type MP Format Metrics Annotation
DocVQA WACV 2021 Industrial Reports English N/A 12,767 50,000 Text N D./P./H. ANLS Human
VisualMRC AAAI 2021 Website English N/A 10,197 30,562 Text N D. BLUE, etc Human
TAT-DQA MM 2022 Financial Reports English 2,758 3,067 16,558 Text/RS-Gen. Y D. EM, F1 Human
RDVQA MM 2022 Data Analysis Report N/A 8,362 8,514 41,378 Text N D. ANLS, ACC Human
CS-DVQA MM 2022 Industry Documents English N/A 600 1,000 Text and Nodes N D./P./H. ANLS Human
PDFVQA-Task A ECML-PKDD 2023 Academic Paper English N/A 12,337 81,085 Num or Yes/No N D. F1 Template
PDFVQA-Task B ECML-PKDD 2023 Academic Paper English N/A 12,337 53,872 Entity N D. F1 Template
PDFVQA-Task C ECML-PKDD 2023 Academic Paper English 1,147 12,337 5,653 Entity Y D. EM Template
MPDocVQA PR 2023 Industrial Reports English 6,000 48,000 46,000 Text Y D./P./H. ANLS Human
DUDE ICCV 2023 Cross-domain English 5,019 28,709 41,541 Text, Yes/No Y D. ANLS Human
MMVQA IJCAI 2024 Academic Paper English 3,146 30,239 262,928 Entity Y D. EM, PM, MR LLM + Human
Table 2. Summary of visually rich document question answering datasets
5.2.1. Single Page VRD-QA Datasets
DocVQA (Mathew et al., 2021) is a pioneering dataset in document-based Visual Question Answering (VQA), sourced from the UCSF Industry Document Library. It comprises 50,000 manually generated questions framed on 12,767 document images, encompassing digital, printed, and handwritten formats. The dataset follows an extractive-style QA format similar to benchmarks like SQuAD (Rajpurkar et al., 2016b) and VQA (Biten et al., 2019b). Evaluation typically involves fine-grained models such as LayoutLM variants (Xu et al., 2020a, b; Huang et al., 2022), LiLT (Wang et al., 2022b), and generative models (Tang et al., 2022; Kim et al., 2022), using metrics like Average Normalized Levenshtein Similarity (ANLS) (Biten et al., 2019a). However, it requires additional processing for coarse-grained models like SelfDoc (Li et al., 2021b). CS-DVQA (Du et al., 2022) builds upon DocVQA by enhancing QA pairs to better reflect real-world requirements. It extracts 600 images from the DocVQA dataset and generates 1,000 QA pairs under human supervision. During question generation, it incorporates common-sense knowledge from real life, expanding answers beyond extractive in-line text to include question-related nodes (Nodes) sourced from ConceptNet (Speer et al., 2017).
VisualMRC (Tanaka et al., 2021) is compiled from website screenshots across 35 domains, carefully selected to exclude pages with handwritten content and to prefer pages containing short text (no more than 2 to 3 paragraphs). Unlike other datasets that might only provide question-answer annotations (Mathew et al., 2021) or automatically acquire document semantic entities (Ding et al., 2023b), VisualMRC includes manually annotated layout structures with fine-grained semantic entity types such as Heading, Paragraph, Subtitle, Picture, and Caption. Question-answer pairs are generated through crowdsourcing. Consequently, VisualMRC is well-suited for evaluating both fine-grained and coarse-grained-based QA frameworks, providing a rich resource for assessing the effectiveness of models in understanding and interpreting detailed document layouts and semantic entities.
PDFVQA-Task A and Task B (Ding et al., 2023b) form part of the first document VQA dataset from PubMed Central, focusing on content and structural understanding. This dataset includes three tasks: two for single-page documents (Tasks A and B) and one for multi-page documents (Task C). Task A evaluates the structural and spatial relationships within document images, with answers being either counts or Yes/No. Task B focuses on extracting document entities based on their logical and spatial configurations. The PDFVQA dataset provides only coarse-grained, entity-level annotations, necessitating further processing for models that require fine-grained analysis. This setup is ideal for testing models’ capabilities in understanding the logical and spatial structures of document images.
5.2.2. Multi-Page VRD-QA Datasets
TAT-DQA (Zhu et al., 2022), an extension of the TAT-QA (Zhu et al., 2021a) dataset, is developed with more complex natural document structures and an expanded set of manually corrected and generated question-answer pairs derived from business financial reports. Unlike other datasets that primarily focus on extractive or simple abstractive answers (such as counting or yes/no), TAT-DQA includes questions requiring arithmetic reasoning, where values must be extracted from tabular data and textual content for discrete calculations. This dataset adopts the evaluation metrics of TAT-QA, including Exact Matching and a numeracy-focused F1 score. These metrics are particularly tailored to assess the accuracy of arithmetic reasoning and data extraction capabilities of the models tested with TAT-DQA.
RDVQA dataset (Wu et al., 2022) compiles a large collection of conversational chats and associated images from an E-commerce platform. It employs standard OCR and Named Entity Recognition (NER) techniques to extract text and redact sensitive information, ensuring privacy protection through masking. The dataset includes question-answer pairs within the images, which are manually verified to confirm image clarity and the presence of at least one question-answer pair per image. Although some documents span multiple pages, this dataset is structured such that it can be processed relatively easily by single-page VRD-QA models.
PDFVQA-Task C (Ding et al., 2023b) is a distinct sub-task within the PDFVQA dataset that expands document VQA to encompass entire long documents, moving beyond the single-page focus of Tasks A and B. In Task C, to answer questions, the model often needs to retrieve information from multiple document entities. Thus, Task C employs Exact Matching for its ground truth annotations. Similar to Tasks A and B, additional processing is required for evaluating models at a fine-grained level.
MP-DocVQA (Tito et al., 2023) extends the original DocVQA (Mathew et al., 2021) dataset to accommodate multi-page document analysis. This version includes adjacent pages from the same documents, expanding the dataset from 12,767 to 64,057 document images. In adapting to a multi-page format, some questions inappropriate were removed. However, it’s important to note that while the dataset allows for questions across multiple pages, the answers remain confined to individual pages; there are no cross-page answers in the MP-DocVQA dataset.
DUDE (Van Landeghem et al., 2023) is the first cross-domain, multi-page document VQA dataset, featuring a diverse collection of documents from various fields such as medical, legal, technical, and financial, and different document types including CVs, reports, and papers. It comprises 5,019 documents, 28,709 document pages, and 41,541 manually annotated questions. Question types vary from extractive in-line text and Yes/No answers to multi-hop reasoning and structural understanding, similar to those in PDFVQA (Ding et al., 2023b). To evaluate model performance, DUDE uses the ANLS metric (Mathew et al., 2021) for assessing answer prediction accuracy. Additionally, it employs two other metrics: Expected Calibration Error (Guo et al., 2017) and Area-Under-Risk-Coverage-Curve (CURC) (Geifman and El-Yaniv, 2017; Jaeger et al., 2022) to gauge the overconfidence and miscalibration in document understanding models. These features make DUDE a comprehensive tool for evaluating cross-domain document understanding models.
MMVQA (Ding et al., 2024a) is a dataset sourced from PubMed Central, designed for the retrieval of multimodal semantic entities from multi-page documents. The questions are generated using ChatGPT (OpenAI, 2023) and subsequently verified manually. Unlike other datasets that focus solely on in-line text or text-dense entities, MMVQA also considers entire tables and figures as potential answers to the given questions. This dataset introduces various evaluation metrics to cater to different application scenarios: Exact Matching and Partial Matching Accuracy assess the precision of responses, while Multi-label Recall evaluates how well the model identifies all relevant answers across the document. This diverse set of metrics makes MMVQA suitable for comprehensive performance evaluation in complex, multimodal document understanding tasks.
5.3. Summary of VRDU Datasets
The common characteristics of KIE and VRD-QA datasets are represented in Table 1 and Table 2. Most existing benchmark datasets for key information extraction are designed for single-page scenarios and predominantly cater to English documents. However, real-world applications often involve more complex multi-page forms. Even forms typically require input from multiple parties and contain multiple languages, such as customs or import/export declaration forms, presenting unique challenges not addressed by current datasets. Similarly, in the domain of VRD-QA, while multi-page VRD-QA has gained increased attention, the integration and exploration of multimodal information remain insufficiently developed. Specifically, structure and relation-aware VRD-QA, which is crucial for interpreting and understanding cross-page relationships, is still a largely unexplored area.
6. Critical Discussion
In Section 3 and Section 3, the models proposed for handling both mono-task and multi-task Visually Rich Document Understanding (VRDU) are introduced. This section aims to provide a summary and analysis of relevant techniques, including their application scenarios, advantages, and disadvantages. Several key topics are covered: Feature Engineering, Cross-Modality Fusion, Model Architecture, and Pre-training Mechanisms. Additionally, with the rapid development of Large Language Models (LLMs), LLM-based frameworks and emerging trends in applying LLMs to VRDU are discussed.
6.1. Feature Representation
6.1.1. Textual Representation
Text in VRDs provides essential semantic context, crucial for understanding content and conducting various downstream tasks. Depending on the information granularity required by the framework and application scenarios, textual representation methods can generally be categorized at the word or entity level.
Figure 4. Fine-grained (word-level) and coarse-grained (entity-level) textual information encoding for VRDU frameworks.
Word-level Representations In Visually Rich Documents (VRDs), text sequences extracted by off-the-shelf OCR tools or PDF parsers (e.g., PDFMiner) can be encoded using word embedding methods such as Word2Vec (Mikolov et al., 2013a) and Glove (Pennington et al., 2014). For more comprehensive textual embeddings, various BERT-style bi-directional pretrained transformer models like BERT (Devlin, 2018) and RoBERTa (Liu et al., 2019b) are employed to generate context-aware word representations. As the visually rich and structurally complex nature of VRDs, layout-aware and visual-integrated fine-grained models have been developed, such as LayoutLM families (Xu et al., 2020a, b; Huang et al., 2022; Xu et al., 2021), LiLT (Wang et al., 2022b). These models generate word representations that integrate multimodal information, combining text, visual cues, and layout structure to achieve SoTA performance on several downstream tasks.
Entity-level Representations To acquire a dense representation of a text sequence for performing entity-level VRDU tasks, various approaches are adopted. These include averaging word embeddings specific to an entity or leveraging the $<CLS>$ token to encapsulate the entire sequence, including averaging the word embeddings belonging to an entity or using $<CLS>$ token to represent an entire sequence. SentenceBERT (Reimers and Gurevych, 2019) is also often adopted to encode text sequences within document entities. However, a standardized approach for acquiring textual representations of entities is yet to be established, necessitating preliminary testing and validation.
6.1.2. Visual Representation
Figure 5. Commonly adopted visual information encoding approaches.
Visual information provides layout, structural insights, and rich contextual clues, making it easier for humans to interpret and prioritize content and resulting in a more comprehensive reading experience. Based on the methods used for encoding visual information, we categorize them into two main types: CNN-based and Vision Transformer-based approaches.
CNN-based Vision Encoding methods involve first acquiring Region of Interest (RoI) bounding boxes and then applying RoI-pooling and RoI-Align on pretrained CNN backbones (e.g., Faster-RCNN or Mask-RCNN) to extract the region features. Many frameworks (Xu et al., 2020a; Li et al., 2021b; Luo et al., 2022; Ding et al., 2024a; Zhai et al., 2023; Ding et al., 2023b; Gu et al., 2021) utilize word or entity-level RoIs to effectively extract visual features of target regions and leverage the implicit knowledge embedded in pretrained backbones. However, acquiring the bounding boxes of words or entities incurs additional costs. Therefore, several frameworks (Xu et al., 2020b, 2021; Bai et al., 2022; Appalaraju et al., 2021) directly use image patch bounding RoIs to extract visual features and learn contextually with other modalities.
After acquiring the visual features, they are typically fed into a transformer framework to fuse multimodal information, which can create significant computational bottlenecks. Additionally, acquiring high-quality RoIs of words or entities requires supervised training. To address these challenges, LayoutLMv3, inspired by ViT (Dosovitskiy et al., 2020), introduces a transformer-only framework. This approach applies a linear layer to project flattened patch pixels, which are then fed into a multimodal transformer to contextually learn with other modalities. This method reduces the number of parameters and simplifies the preprocessing steps, making it more efficient and adopted by many recent frameworks (Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024). However, this encoding method cannot take advantage of implicit knowledge in pre-trained frameworks and typically requires extensive pre-trained.
6.1.3. Layout Representation
Figure 6. Commonly adopted layout information encoding approaches.
Layout information is crucial for understanding document elements’ spatial arrangement, including words and entities. Enhance document representation by clarifying the spatial relationships between these elements, thereby aiding in the comprehension of the overall document structure. The coordinates of the bounding box (bbox) of the document elements serve as initial layout information. This layout information can then be encoded using methods such as positional encoding, linear projection, and spatial-aware attention bias.
2D positional encoding, first introduced by LayoutLM (Xu et al., 2020a), and widely adopted by many visually rich document understanding (VRDU) models, allows the model to be aware of the relative spatial positions within a document. In this approach, document elements are normalized and discretized into integer ranges, and two separate embedding layers are used to encode the x and y coordinates, respectively. Despite its widespread use in models like (Xu et al., 2020a, b, 2021; Huang et al., 2022; Tu et al., 2023; Wang et al., 2022b; Bai et al., 2022; Liao et al., 2023)this method encodes x and y coordinates individually, making it challenging to represent continuous 2D space and capture special correlations between document elements. Some models (Ding et al., 2023a; Wang et al., 2020b) that follow the approach of LXMERT (Tan and Bansal, 2019) utilize linear projection to update the x and y coordinates of the normalized bounding box coordinates simultaneously. To address the limitations of absolute positional encoding and incorporate relative positional correlations, other models introduce spatial-aware attention mechanisms (Hong et al., 2021; Appalaraju et al., 2021; Wang et al., 2022b). These mechanisms enable vanilla self-attention to learn spatial dependencies effectively.
6.2. Multi-modality Fusion
Figure 7. Commonly adopted multi-modality fusion methods.
After acquiring multimodal representations, it is important to explore effective fusion methods to integrate textual, visual, and layout information. This integration improves document understanding and boosts performance on downstream tasks. The straightforward integration methods, as shown in Figure 7, include the additive and concatenation of the feature vectors. For example, additive integration sums layout information with corresponding textual or visual token representations (Xu et al., 2020a, b; Huang et al., 2022; Li et al., 2021b), while concatenation merges visual and textual features of document entities (e.g. tables) (Ding et al., 2024a, 2023b). However, these methods require one-to-one correlations and alternative approaches are needed when such correlations are not available. Consequently, self-attention and cross-attention mechanisms are widely adopted to enhance each modality by learning inter- and intramodality contexts. These mechanisms are commonly used in frameworks (Xu et al., 2020b; Huang et al., 2022; Tu et al., 2023; Bai et al., 2022; Xu et al., 2021) that integrate patch-level visual embeddings with textual features for contextual learning. Novel self-attention (Appalaraju et al., 2021; Wang et al., 2022b) and cross-attention (Gu et al., 2021; Li et al., 2021b; Zhai et al., 2023) methods have been proposed to fuse multimodal information more effectively. Apart from model-based fusion approaches, self-supervised and joint learning tasks are also effective for integrating multi-aspect features. Self-supervised pretraining tasks such as Masked Visual-Language Modeling (Wang et al., 2022b; Xu et al., 2020a, b; Li et al., 2021a; Huang et al., 2022; Zhai et al., 2023), Text-Image Alignment (Appalaraju et al., 2021), Text-Layout Pairing (Wang et al., 2022b), and Text-Image Matching (Xu et al., 2020b, 2021) can significantly enhance multimodal information fusion. These methods require large-scale pretraining to learn cross-modality semantic correlations. Joint learning methods, often used in OCR-free frameworks (Wang et al., 2021a; Kim et al., 2022; Davis et al., 2022; Yu et al., 2022), design auxiliary text detection or recognition tasks to fuse textual and visual information. This approach reduces pre-processing during inference and addresses mis-ordering sequence issues. However, these methods generally underperform compared to OCR-dependent models and involve additional training costs.
6.3. Model Architecture
6.3.1. Transformer in VRDU
Referring to the models introduced in Section 3 and Section 3, transformers have become extensively utilized in VRDU tasks, attaining state-of-the-art performance due to several key advantages. Firstly, the attention mechanism effectively captures long-range dependencies within the multimodal information, including text, vision, and layout. Furthermore, the inherent scalability of transformers improves self-supervised learning on large-scale datasets (e.g. IIT-CDIP (Lewis et al., 2006)), allowing them to handle diverse types and formats of documents, capturing more intricate and comprehensive features of the documents. Based on the transformer architecture used, VRDU models can be divided into two categories: encoder-only and encoder-decoder-based models. The first encoder-only model, LayoutLM (Xu et al., 2020a), was inspired by BERT (Devlin, 2018) and uses various pretraining tasks to allow the bidirectional transformer encoder to capture more textual and layout information. Following LayoutLM, more pretrained VRDU models with encoder-only, layout-aware (Wang et al., 2022b; Hong et al., 2021; Chen et al., 2022; Wang et al., 2022b; Tu et al., 2023; Li et al., 2021a) or visual integrated (Xu et al., 2020b, 2021; Appalaraju et al., 2021; Huang et al., 2022) have been proposed. These models are pretrained on various tasks to enhance their understanding of document structures. Encoder-only models demonstrate remarkable performance in sequence tagging and document classification tasks. However, they face challenges related to heavy annotation requirements and low readability, and they struggle with generative tasks such as abstractive question answering. Additionally, OCR errors can complicate the extraction of accurate information from input text. Furthermore, the fixed maximum input length of encoder-only frameworks limits their ability to handle long document inputs effectively. To overcome the limitations of encoder-only frameworks in generative QA and KIE, several encoder-decoder models (Powalski et al., 2021; Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024) have been developed, but they still depend on costly OCR tools which can introduce errors affecting performance. OCR-free frameworks (Kim et al., 2022; Davis et al., 2022; Cao et al., 2023a; Yu et al., 2022) address this issue by using vision encoders and text decoders for end-to-end processing. For long documents, T5-based encoder-decoder models (Raffel et al., 2020) have been proposed to effectively handle multipage contexts.
6.3.2. CNNs in VRDU
In VRDU, CNNs are employed as the core framework for extracting feature maps from document images (Yang et al., 2017; Palm et al., 2019) and character grids (Katti et al., 2018), benefiting from their strong local feature extraction capabilities. Joint-learning frameworks (Wang et al., 2021a, a) use CNNs as a backbone to integrate OCR and KIE tasks, combining visual and textual information through auxiliary tasks. Some OCR-free pretrained frameworks (Davis et al., 2022; Yu et al., 2022) also utilize CNNs as vision encoders to extract visual feature maps. However, CNNs struggle to capture long-range dependencies due to their localized receptive fields. To address this, StrucTexTv2 (Yu et al., 2022) combines CNN-extracted feature maps with a transformer to capture global contextual information. Additionally, CNNs are commonly used to extract visual features from regions of interest (RoI) with RoI Align(Xu et al., 2020a, b, 2021; Huang et al., 2022; Tu et al., 2023; Wang et al., 2022b; Bai et al., 2022; Liao et al., 2023). Although CNNs can accurately capture region-specific visual cues and leverage pre-trained knowledge from general domains, they require extra processing to obtain RoI bounding boxes, unlike vision transformers (Huang et al., 2022; Tang et al., 2022; Appalaraju et al., 2024; Mao et al., 2024), which operate directly on patch pixel values.
6.3.3. Graphs in VRDU
VRDs feature complex spatial and logical structures. The spatial structure shows the layout and positional relationships, such as a Title above a Paragraph and a Caption near a Figure or Table. The logical structure denotes semantic and hierarchical connections, like a Title being the parent of a Paragraph and a Caption describing a related Table or Figure. Graph-based frameworks explicitly encode these relationships using node and edge representations; therefore, Graph Neural Networks (GNNs) are widely used in VRDU models (Lee et al., 2022b; Shi et al., 2023; Luo et al., 2022; Zhang et al., 2021) to encode the spatial and logical representations. Although GNNs effectively capture domain-specific knowledge, they struggle with scalability and general domain knowledge pretraining. To address this, some frameworks (Li et al., 2023; Zhang et al., 2022) use attention masks or biases to mimic relationships between document elements, blending attention mechanisms with explicit relational modelling.
6.4. Pretraining Mechanisms
By performing various pretraining tasks, a model can enhance its generalization ability through extensive datasets and prior training. Inspired by advances in pre-training language (Devlin, 2018; Brown et al., 2020) and vision models (He et al., 2016; Dosovitskiy et al., 2020), numerous pretraining tasks for VRDU have been developed that are typically trained in large-scale document collections. This section will summarize the pretraining techniques and datasets commonly used for VRDU pretraining.
6.4.1. Pretraining Tasks
Based on the purpose and pretraining targets, the pretraining methods can be categorised into Masked Information Modeling, Cross-modality Learning, Mono-modality Augmentation, and Contrastive Learning.
Masked Information Modelling (MIM) is first introduced by Masked Language Modelling in BERT (Devlin, 2018), which randomly masks 15% workpiece tokens and requires the model to predict the masked tokens. Some models use multi-source inputs, such as XDoc (Chen et al., 2022) and MarkupLM (Li et al., 2022a), which directly adopts Masked Language Modeling as a pretraining task on plain text or markdown text subsets. Some methods optimise MLM by changing masking wordpieces to whole words (Whole Word Masking by LayoutMask (Tu et al., 2023)) or all tokens belong to one randomly generated text block, named Area-Masked Language Modelling (Hong et al., 2021). Similar strategies can also be applied to mask entity-level textual representation, e.g. Masked Sentence Modelling (Gu et al., 2021). Those language-targeted masked information modelling to improve the language understanding ability of VRDU models. Except for language-focused masking strategies, vision (Gu et al., 2021; Yu et al., 2022) and layout-focused (Tu et al., 2023; Wang et al., 2022b) strategies are also adopted. Moreover, masked information modelling is an effective method to boost cross-modality understanding. LayoutLM (Xu et al., 2020a) introduces a Masked Visual-Language Modelling which allows using kept visual/layout information and contextual text content to predict the masked word-pieces, adopted by many VRDU pretrained models (Wang et al., 2022b; Xu et al., 2020b; Li et al., 2021a; Huang et al., 2022; Zhai et al., 2023; Appalaraju et al., 2021). Similarly, some visual token masked models leverage multimodal information to reconstruct the view tokens, e.g. Learn to Reconstruct (Appalaraju et al., 2021), Masked Image Modeling (Huang et al., 2022). Additionally, some models mask multimodal features simultaneously to conduct a cross-modality masking (Li et al., 2021b; Liao et al., 2023). Masking Information Modelling may have limitations on bias in masking strategies, thus some frameworks (Tu et al., 2023) may try different masking ratios or strategies to improve the training effectiveness. Additionally, other common concerns about MIM also include training efficiency and lack of structural information. Thus, other pretraining methods are introduced to mitigate the limitations.
Cross-modality Aligning Although some Masked Information Modelling methods could effectively boost the cross-modality understanding, implicit contextual learning is limited to capturing explicit alignment between different modalities. Thus, few cross-modality aligning methods are introduced to enhance the modality interaction. To enhance the text-image interactive learning, Text-Image Alignment is adopted (Xu et al., 2020b, 2021) which covers the image region of token lines to predict whether the image region of the target token line is covered or not. LayoutLMv3 (Huang et al., 2022) expands from covering the image region only to image/text tokens to further enhance interactive learning. Vision-language alignment (Gu et al., 2021) and Text-Image Matching (Xu et al., 2020b, 2021) target to predict whether image-text features belong to the same region or not. DocFormer (Appalaraju et al., 2021) tends to predict the text content of the paired images.
Other Pretraining Techniques: Some pretraining tasks are introduced to further enhance the understanding of specific modalities. To enhance layout information understanding, StructuralLM (Li et al., 2021a) and WOKONG-READER (Bai et al., 2022) introduce Cell Position Classification and Textline Grid Matching to predict which located grids are of each cell or textline. Fast-StrucText (Zhai et al., 2023) introduces a graph-based token relation method to predict a spatial correlation between token pairs. MarkupLM (Li et al., 2022a) leverage the benefits from markup files to predict the logical parent-child relation between nodes by introducing a Node Relation Prediction. Contrastive learning based strategies are adopted to conduct single (Gu et al., 2021) or cross-modality contrastive (Textline-Region Contrastive Learning) (Bai et al., 2022) learning.
6.4.2. Pretraining Datasets
To perform the aforementioned tasks, large-scale document collections are essential for conducting self-supervised learning. Different pretraining datasets are adopted by various models. The most widely used pretraining dataset is the IIT-CDIP Test Collection 1.0 (Lewis et al., 2006), which contains more than 6 million documents with over 11 million scanned document images. Since it contains a cross-domain and large number of unannotated documents, it is used by the majority of models (Li et al., 2022b; Hong et al., 2021; Kim et al., 2022; Xu et al., 2020a, b; Huang et al., 2022; Xu et al., 2021; Bai et al., 2022; Appalaraju et al., 2021; Liao et al., 2023; Wang et al., 2022b; Gu et al., 2021; Zhai et al., 2023; Li et al., 2021a; Luo et al., 2023; Chen et al., 2022; Appalaraju et al., 2024). As the original IIT-CDIP dataset provides the text content without layout information, off-the-shelf OCR tools are normally used to acquire the bounding box information of each document. Some models (Li et al., 2021c, b; Wang et al., 2022a) use relatively smaller pretraining datasets like RVL-CDIP (Harley et al., 2015), which contains 400,000 evenly distributed documents in 16 types, to reduce the cost of pretraining. The multi-source model XDoc (Chen et al., 2022) also leverages many plain text corpora for pretraining, such as BookCORPUS, CC-NEWS, OPENWEBTEXT, STORIES and HTML-sourced CommonCrawl datasets. To address multilingual scenarios, both LiLT (Wang et al., 2022b) and LayoutXLM (Xu et al., 2021) follow the principles and policies of Common Crawl to gather large amounts of multilingual digitally-born PDF documents.
7. Conclusion
This paper comprehensively reviews deep learning-based models for visually rich document content understanding, encompassing both mono-task frameworks designed for specific VRDU downstream tasks and multi-task frameworks that support multiple VRDU downstream tasks. Beyond introducing the novelties of each model, the limitations of these frameworks are summarized at the end of each section, offering a thorough trend analysis. Additionally, this paper summarizes existing VRD content understanding datasets, pointing out future trends and demands for VRDU. To provide a systematic review, we critically discuss various techniques, highlighting their strengths and limitations. We believe that this survey offers a comprehensive overview of the development of visually rich document content understanding, catering to the needs of both the academic and industrial sectors.